### v1

# ORMNet: Object-centric Relationship Modeling for Egocentric Hand-object Segmentation

Abstract
Egocentric hand-object segmentation (EgoHOS) is a brand-new task aiming at segmenting the hands and interacting objects in the egocentric image. Although significant advancements have been achieved by current methods, establishing an end-to-end model with high accuracy remains an unresolved challenge. Moreover, existing methods lack explicit modeling of the relationships between hands and objects as well as objects and objects, thereby disregarding critical information on hand-object interaction and introducing confusion into algorithms, ultimately leading to a reduction in segmentation performance. To address the limitations of existing methods, this paper proposes a novel end-to-end Object-centric Relationship Modeling Network (ORMNet) for EgoHOS. Specifically, based on a single-encoder and multi-decoder framework, we design the Hand-Object Relation (HOR) module to leverage hand-guided attention to capture the correlation between hands and objects and facilitate their representations. Moreover, based on the observed interrelationships between diverse categories of objects, we introduce the Object Relation Decoupling (ORD) strategy. This strategy allows the decoupling of the two-hand object during training, thereby alleviating the ambiguity of the network. Experimental results on three datasets show that the proposed ORMNet has notably exceptional segmentation performance with robust generalization capabilities.
Index Terms:
Egocentric view, hand-object interaction segmentation, attention, relationship modeling.
I Introduction
The advancements in edge computing and computer vision have engendered a growing interest in Head-Mounted Devices (HMD) such as the Apple Vision Pro[1] and Microsoft Hololens 2[2]. Consequently, there is a rising trend in capturing abundant first-person view (FPV) or egocentric images and videos. Compared with data captured from a third-person view (TPV) or an exocentric perspective[3], egocentric data displays the actions and intentions of the HMD wearer more clearly and intuitively, which is more conducive to analyzing human behavior and understanding how humans interact with the environment. Notably, questions concerning the activities of individuals, such as “What is this person doing?” can be more comprehensively addressed through the parsing of egocentric visual data. It is obvious that the hands of HMD wearers and the objects they are interacting with serve as pivotal cues in this analysis.
Therefore, to delve deeply into human behavior encoded within egocentric data at a fine-grained level, Zhang et al.[4] first proposed the Egocentric Hand-Object Segmentation (EgoHOS) task, which aims to perform pixel-level segmentation of the hand and the objects that interact with the hand in egocentric images. This novel segmentation task enables a precise response to the query, “Where are the left and right hands, and what objects are they interacting with?” which can function as an essential algorithm serving as the foundation for various subsequent applications, such as augmented reality/virtual reality (AR/VR)[5, 6], medical assistant[7, 8], robotics[9], embodied AI[10, 11], and metaverse[12].
Figure 1: Structural differences between the previous (a) parallel method, (b) sequential method, and (c) our approach. Compared with previous parallel and sequential methods, the proposed ORMNet can achieve higher segmentation performance with higher efficiency by integrating relationship modeling between “hands and objects (H-O)” as well as “objects and objects (O-O).” The $N,N_{h},$ and $N_{o}$ mean the number of overall categories, the number of hand categories, and the number of object categories, and $N=N_{h}+N_{o}$.
The previous approaches for EgoHOS can be categorized into parallel and sequential methods. Specifically, parallel methods[4, 13, 14, 15, 16] are straightforward, which employ an end-to-end encoder-decoder framework to extract features and predict the mask of all categories simultaneously, as shown in Fig. 1(a). In contrast to parallel methods, sequential approaches[4, 17] employ a step-by-step prediction scheme for the hand and interacting objects, i.e., the input of the latter step to predict the objects’ masks necessitates the previous step’s output for predicting the hand, as depicted in Fig. 1(b).
Although existing methods have demonstrated a satisfied EgoHOS capability, they are accompanied by many drawbacks, as shown in the table in Fig. 1. First, similar to conventional encoder-decoder semantic segmentation networks[18, 19], the intuitive parallel approaches[4] regard hands and objects as independent categories and predict segmentation results for them in an end-to-end manner. While this type of technique demonstrates high efficiency, it fails to capture the relationship between “hands and objects” or among different objects, consequently leading to performance deficiencies. Second, in contrast, the sequential methods[4] specially built for the EgoHOS task demonstrate enhanced performance by incorporating the relationships between hands and interacting objects. However, step-wise training and inference elongate the overall time consumption, thereby reducing the method’s efficiency and constraining its feasibility for real-time applications. Third, the prediction of diverse interacting objects is required in the EgoHOS task, i.e., left-hand objects (objects interact with the left hand), right-hand objects (objects interact with the right hand), and two-hand objects (objects interact with both hands at the same time). This paper observes the coupling relationship between these three categories of objects, i.e., the two-hand object can be regarded as both a left-hand object and a right-hand object. Therefore, it is easy to predict a two-hand object as the left-hand object when the left interaction is more obvious than the right interaction, and the same applies to the right-hand object. Nevertheless, prevailing methods overlook the ambiguity of such relationships, which introduces confusion during network training.
To improve efficiency and effectiveness simultaneously, this paper proposes a novel end-to-end Object-centric Relationship Modeling Network (ORMNet) to model the relationship between “hands and objects” as well as “objects and objects,” as shown in Fig. 1(c). On the one hand, we propose the Hand-Object Relation (HOR) module to establish the correlation between the features of the hand and interacting objects. The core underlying concept involves leveraging hand features as prior knowledge and employing hand-guided attention to extract more representative object features. On the other hand, we introduce a new Object Relation Decoupling (ORD) strategy to model the relationship among diverse object categories. Drawing from the insights gained from analyzing two-hand objects, this strategy restricts the prediction to left- and right-hand objects exclusively during the training phase. Then, two-hand objects are generated through a simple yet efficient intersection operation during inference. This strategy can reduce the confusion of the algorithm about the object categories during learning, enabling the network to concentrate more precisely on the interacting objects. The contributions of this paper are as follows:
  • This paper introduces a novel end-to-end ORMNet for the EgoHOS task that achieves state-of-the-art segmentation performance while being highly efficient.
  • To take full advantage of the prominence of hand features, the proposed HOR module models the relationship between hands and objects, helping the network extract more comprehensive hand and object features.
  • Based on the relation of various object categories, the simple yet powerful ORD strategy is proposed to decouple the two-hand objects during training, reducing the ambiguity of the network and further separately highlighting the objects contacted with left or right hands.
  • The experimental results on two datasets demonstrate that the proposed ORMNet achieves exceptional performance and manifests robust generalization capabilities.
II Related Works
II-A Image Segmentation
Image segmentation is a basic task of computer vision, which aims to divide the image into several non-overlapping regions based on certain principles, such as semantic information. Semantic segmentation[20, 21, 22], instance segmentation[23, 24, 25], and panoptic segmentation[26, 27, 28] are three commonly studied segmentation sub-tasks. Image segmentation methods can be divided into traditional methods, convolution neural network (CNN)-based methods, and transformer-based methods in chronological order. Although traditional methods can perform image segmentation with notable interpretability, further refinement is needed to enhance the quality of the segmentation outcomes. The fully connected neural network (FCN)[19] proposed by Jonathan et al. is the cornerstone of CNN segmentation methods. After that, there have been numerous advancements[29] in CNN-based image segmentation methods, such as U-Net[18], Deeplab-series[30, 31], SegNet[32], PSPNet[33], etc.
The development of transformer-based methods for image segmentation is a relatively recent but rapidly evolving field. Dosovitskiy et al. introduced the Vision Transformer (ViT)[34] in 2020, proving that the transformer could be applied directly to images by treating the image as a sequence of patches. Following the success of ViT, several transformer-based methods[35] specially designed for image segmentation have been proposed. Zheng et al. developed the SETR[36] to adopt ViT as a backbone to extract features to eliminate the reliance on FCN, achieving impressive performance. Similarly, Strudel et al. proposed a pure end-to-end Transformer approach - Segmenter[37], which utilized the ViT encoder for capturing global context and a simple point-wise linear decoder to generate precise segmentation masks. Meanwhile, Xie et al. presented the Segformer[13] to integrate hierarchical Transformer encoder with lightweight multi-layer perceptron (MLP) decoders, avoiding complex designs in segmentation methods. Unlike previous work, Cheng et al. proposed the MaskFormer[38], which abandoned the per-pixel classification formulation and instead predicted a series of binary masks associated with class labels. Following that, Cheng et al. introduced the Mask2Former[39], which incorporated masked attention to perform universal (i.e., instance, semantic, and panoptic) segmentation, achieving notable performance. Then, Jain et al. proposed OneFormer[14] to perform universal segmentation tasks with transformers and task-guided queries, which achieved state-of-the-art performance in various segmentation benchmarks. Recently, Kirillov et al. proposed a new impressive work called Segment Anything (SAM)[40]. They built a foundation model for interactive segmentation by training with over 1 billion masks on 11M images.
Our paper utilizes the Swin transformer[41] to extract the features of the input egocentric image and to up-sample the features to the input size[15]. Although using more recent methods may improve performance, this is not the main focus of this paper.
II-B Egocentric Hand-object Segmentation
Zhang et al.[4] recently introduced a novel task known as Egocentric Hand-object Segmentation (EgoHOS), which combines image segmentation and Human-Object Interaction detection (HOI-det). This task redefines the focus of HOI-det from the entire human body to specific, fine-grained parts, i.e., human hands. Its objective is to segment the left hand, right hand, and the objects that interact with the left and right hands in input RGB egocentric images. Concurrently, Zhang et al. [4] introduced the innovative concept of the contact boundary (CB), representing the region where the hands and the objects engage in interaction. This concept explicitly promotes the network focus on the interaction between hands and objects. Furthermore, based on the proposed task and concept, the authors also proposed different types of methods in their work, including parallel methods and sequential methods.
Parallel method. A more direct and naive method is to use an encoder and a decoder to segment and predict hands and objects and use different indexes to imply different categories[4, 13, 14, 15, 16], as shown in Fig. 1(a). Although this parallel method completes the prediction of this task simply and directly, its segmentation accuracy is relatively low and still needs to be improved.
Sequential method. Since parallel methods lacked targeted design in learning hand-object interaction, Zhang et al. proposed a sequential method[4], as shown in Fig. 1(b). Taking the egocentric image as input, this step-by-step sequential method uses an encoder-decoder first to segment the left and right hands. After that, the segmented left- and right-hand masks are concatenated to the input image to be used to perform the mask prediction of contact boundary and interacting objects. Although the sequential method achieved higher segmentation performance than the parallel method, it lacks efficiency due to the step-by-step process.
In addition, both parallel and sequential methods inadequately explore the relationship between hands and objects and ignore the relationship between objects and objects. Unlike the previous parallel and sequential methods, the proposed ORMNet can achieve end-to-end training and inference with high accuracy due to the relationship modeling between “hands and objects” as well as “objects and objects.”
II-C Attention Mechanism in deep learning
Models using the attention mechanism have exploded in recent years, confirming the effectiveness of the attention mechanism for NLP[42, 43] and computer vision[44, 45] tasks. The primary objective of attention in vision is to mimic the selective function of the human visual cognitive system towards input information, i.e., concentrating on the primary and crucial features of the input visual data. The majority of attention modules utilize distinct methodologies to generate an attention map for feature weighting, consequently adjusting the significance of individual components within the features.
In the early stage, most attention modules are integrated into the CNN-based models. For example, Hu et al. proposed the SENet[46] equipped with channel-wise attention, which used global average pooling to generate the attention weight among channel axial. After that, Woo et al. proposed CAM [47] to improve channel attention by aggregating the descriptors from max-pooling and average-pooling. There are many other methods intended to improve channel-wise attention, e.g., SOCA[48], HA[49], etc. Unlike channel attention, spatial attention is proposed to generate weights for spatial patches of feature maps. For example, Woo et al. proposed SAM [47] to use a convolutional layer to obtain the attention weights.
The development of self-attention and multi-head self-attention (MHSA) mechanisms in transformer-based models has also been widely used in various methods in recent years. Some researchers also make an effort to extend self-attention to cross-attention[50, 51], which allows the model to capture relationships between different modalities of data[52, 53]. The proposed ORMNet utilizes hand-guided cross-attention to model the relationship between the features of hands and objects, allowing more comprehensive feature extraction.
Figure 2: Overall diagram of the proposed ORMNet, which consists of an encoder, bottleneck, multiple decoder branches, Hand-Object Relation (HOR) module, and Object Relation Decoupling (ORD) strategy. The algorithm is utilized to predict the hands, right-hand objects (RO), left-hand objects (LO), and two-hand objects (TO). The ORD strategy processes the initial (Ini.) ground truth (GT) to generate new supervision during data preparation, and it also processes the predictions (Pred.) of diverse categories of interacting objects during inference.
III Methodology
This section introduces the proposed method in detail from top to bottom, including task description (in Sec. III-A), overall architecture (in Sec. III-B), Object Relation Decoupling Strategy (in Sec. III-E), Hand-Object Relation module (in Sec. III-D), and training and inference (in Sec. III-F).
III-A Task Description
The EgoHOS dataset [4] includes the label of the hands, $1^{st}$-order interacting objects, and $2^{nd}$-order interacting objects. In this paper, we focus on the segmentation of hands and $1^{st}$-order interacting objects, i.e., objects that are directly contacted by hands. Given an egocentric image I, the goal of the EgoHOS task is to predict the masks of the following entities: left hand $\textbf{M}_{H}^{L}$, right hand $\textbf{M}_{H}^{R}$, left-hand objects $\textbf{M}_{O}^{L}$ (objects directly interacting with the left hand), right-hand objects $\textbf{M}_{O}^{R}$ (objects directly interacting with the right hand), and two-hand objects $\textbf{M}_{O}^{T}$ (objects interacting with both hands simultaneously).
III-B Overall Architecture
The Overall diagram of the proposed ORMNet is shown in Fig. 2. The model consists of an encoder, a bottleneck, and a decoder with multiple branches. The encoder is used to extract distinguishing features of the input egocentric image, the bottleneck is used to enhance the extracted global features, and different branches of the decoder are used to predict the final segmentation masks for diverse categories. Following the previous work[4], ORMNet also contains the contact boundary (CB) decoder branch to predict the mask of the contact boundary. For simplification, we remove the CB decoder branch in Fig. 2.
III-C Encoder and Bottleneck
Taking an egocentric image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ as input, the first step of ORMNet is to use the image encoder to extract the multi-scale features $\textbf{F}_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times C_{i}}$ by a series of down-sampling by a factor of 2, where $H_{i}$, $W_{i}$, and $C_{i}$ are the height, width, and channel of the $i^{th}$ stage feature map, respectively. The image encoder used in practice is the Swin Transformer encoder[41]. The extraction process of generating multi-scale features in the encoder can be expressed as follows:
$\mathbf{F}_{i+1}=f_{enc}^{i+1}(\mathbf{F_{i}}|W_{enc}^{i+1}),i\in\left\{0,1,2\right\},$ (1)
$\mathbf{F_{0}}=f_{enc}^{0}(\mathbf{I}|W_{enc}^{0}),$ (2)
where the $f_{enc}^{i}(\cdot)$ denotes the encoder stage with the down-sampling for generating the $i^{th}$ stage feature map, and the $W_{enc}^{i}$ represents corresponding parameters.
After generating the multi-scale features, the final feature map with the smallest height and width and the largest channel is sent into the bottleneck to enhance the representation of the encoded feature. Following the work[15], the bottleneck is composed of a patch merging layer and two Swin Transformer blocks , which is denoted as follows:
$\mathbf{F}_{ehc}=f_{bot}(\mathbf{F}_{last}|W_{bot}),$ (3)
where the $\mathbf{F}_{last}$ and $\mathbf{F}_{ehc}$ represent the final encoded feature map and enhanced final feature map, and the $f_{bot}(\cdot)$ denotes the bottleneck network with the parameter of $W_{bot}$. The obtained enhanced memory feature $\mathbf{F}_{ehc}$ and multi-scale features $\{\mathbf{F}_{i}|i=0,1,2\}$ are then passed into separate decoder branches to predict the hands and interacting objects.
III-D Hand-Object Relation Module
Figure 3: Detailed architecture of the proposed HOR module. Taking the features of hands and objects as input, this module accomplishes hand-guided attention to model the relationship between the hands and interacting objects, promoting the object feature learning of the network.
Among the egocentric visual information, the hand information of the HMD wearer is the most prominent and representative. Moreover, the interacted objects can only be identified based on whether they are in contact with the hand rather than appearance and texture cues. Therefore, using hand features to enhance the extraction and expression of object features is beneficial to the EgoHOS task, which has been ignored in previous work. This paper proposes the HOR module to model the correlation between hand features and object features in interaction. Specifically, this module uses representative hand features as priors and guidance and utilizes the hand-guided attention mechanism to guide the learning of object features, as shown in Fig. 3.
Taking the hand features $\textbf{F}_{hand}\in\mathbb{R}^{\hat{C}\times\hat{H}\times\hat{W}}$ and the object features $\textbf{F}_{obj}\in\mathbb{R}^{\hat{C}\times\hat{H}\times\hat{W}}$ generated by decoder branches as the input, this module aims to use hand features to enhance the object features to focus more precisely on the interacting region, where $\hat{C}$, $\hat{H}$, and $\hat{W}$ are the channel, height, and width of the feature map. First, two stacked convolution operations are used after the LayerNorm layer to generate the query, key, and value. The query $Q\in\mathbb{R}^{\hat{C}\times\hat{H}\hat{W}}$ is generated from hand features, and the key $K\in\mathbb{R}^{\hat{C}\times\hat{H}\hat{W}}$ and value $V\in\mathbb{R}^{\hat{C}\times\hat{H}\hat{W}}$ are generated by object features, which can be denoted as follows,
$Q=f_{conv}(f_{conv}(f_{ln}(\textbf{F}_{hand}))),$ (4)
$K,V=f_{conv}(f_{conv}(f_{ln}(\textbf{F}_{obj}))).\vspace{-0.3cm}$ (5)
After that, the query and key are used to perform the hand-guided attention operation, i.e.,
$\textbf{A}=\delta(Matmul(Norm(Q),Norm(K^{T}))),$ (6)
$\textbf{F}_{o}=Conv(Matmul(\textbf{A},V)),$ (7)
where the $\textbf{F}_{o}$ is the output after convolution, the $Matnul$ means the matrix multiplication, and the $\delta(\cdot)$ represents the softmax function. Then, the input hand features are added to the $\textbf{F}_{o}$ to generate the guided features. Finally, the addition between guided features and original object features is performed to enhance the representation of the object features, which can be described as follows,
$\textbf{F}_{guide}=\textbf{F}_{o}+\textbf{F}_{hand},\quad\hat{\textbf{F}}_{obj}=\textbf{F}_{obj}+\textbf{F}_{guide},$ (8)
where the $\hat{\textbf{F}}_{obj}$ is the output enhanced object features.
III-E Object Relation Decoupling Strategy
Based on the proposed HOR module, a naive way is to use one decoder branch to predict hands and another decoder branch to predict interacting objects. Then, the HOR module can establish the relation between hands and objects, as shown in Sec. III-D.
However, this paper argues that in this way, the three different types of objects, i.e., left-hand objects, right-hand objects, and two-hand objects, are coupled, which will cause confusion and ambiguity when training the network. In particular, if the network is trained to predict left-hand objects and right-hand objects based on whether they are in contact with the left and right hands, then two-hand objects that are in contact with both hands simultaneously are likely to be confused with either left-hand objects or right-hand objects.
Therefore, different from previous work, our method proposes the ORD strategy to decouple the three kinds of objects according to their physical meanings. Specifically, we consider two-hand objects to be both left-hand and right-hand objects so that two-hand objects can be represented by the intersection of the other two categories. This approach establishes a relationship between two-hand objects and the other two categories that are originally unrelated, enabling the network to concentrate solely on the hand-object interaction without classifying the category.
In practice, the ORD process can be divided into data preparation, training, and inference. First, in the phase of data preparation, we consider the ground truth (GT) of two-hand objects in the dataset as both left-hand objects and right-hand objects, so we generate the new GT of left- and right-hand objects by union operation, as shown in Fig. 2 (left). For example, if the value in the two-hand object GT $\textbf{G}_{to}$ at the position $(i,j)$ is 1, then the value of the left-hand object GT $\textbf{G}_{lo}$ and right-hand object GT $\textbf{G}_{ro}$ at $(i,j)$ will be set to 1. The process is shown below,
$\hat{\textbf{G}}_{lo}=\textbf{G}_{lo}\bigcup\textbf{G}_{to},\quad\hat{\textbf{G}}_{ro}=\textbf{G}_{ro}\bigcup\textbf{G}_{to},$ (9)
where the $\bigcup$ means the union operation, the $\hat{\textbf{G}}_{lo}$ and $\hat{\textbf{G}}_{ro}$ are the generated GT labels for left-hand objects and right-hand objects.
Then, in the training process, we only use $\hat{\textbf{G}}_{lo}$ and $\hat{\textbf{G}}_{ro}$ to supervise the two object prediction branches, as shown in Fig. 2, which can be expressed as follows,
$\hat{\textbf{M}}_{O}^{L}=f_{dec}^{lo}(\textbf{F}_{ehc},\textbf{F},W_{dec}^{lo}),\textbf{F}=\{\textbf{F}_{i}|i=0,1,2\},$ (10)
$\hat{\textbf{M}}_{O}^{R}=f_{dec}^{ro}(\textbf{F}_{ehc},\textbf{F},W_{dec}^{ro}),\textbf{F}=\{\textbf{F}_{i}|i=0,1,2\},$ (11)
$\mathscr{L}_{lo}=l_{ce}(\hat{\textbf{M}}_{O}^{L},\hat{\textbf{G}}_{lo}),\quad\mathscr{L}_{ro}=l_{ce}(\hat{\textbf{M}}_{O}^{R},\hat{\textbf{G}}_{ro})),$ (12)
where the $\hat{\textbf{M}}_{O}^{L}$ and $\hat{\textbf{M}}_{O}^{R}$ are the predicted left-hand objects and right-hand objects. The $f_{dec}^{lo}(\cdot)$ and $f_{dec}^{ro}(\cdot)$ represent the decoder branch of left-hand objects and right-hand objects with parameters of $W_{dec}^{lo}$ and $W_{dec}^{ro}$. And the $\mathscr{L}_{lo}$ and $\mathscr{L}_{ro}$ denote the loss function of left-hand objects and right-hand objects, which we use the cross entropy $l_{ce}(\cdot)$ in practice.
In inference, based on the predicted left- and right-hand objects $\hat{\textbf{M}}_{O}^{L}$ and $\hat{\textbf{M}}_{O}^{R}$, the two-hand objects are generated by a simple but effective operation, i.e., intersection, as shown in Fig. 2 (right) and follows,
$\textbf{M}_{O}^{T}=\hat{\textbf{M}}_{O}^{L}\bigcap\hat{\textbf{M}}_{O}^{R},$ (13)
$\textbf{M}_{O}^{L}=\hat{\textbf{M}}_{O}^{L}-\textbf{M}_{O}^{T},\quad\textbf{M}_{O}^{R}=\hat{\textbf{M}}_{O}^{R}-\textbf{M}_{O}^{T},$
where the $\bigcap$ means the intersection operation, and the $\textbf{M}_{O}^{L}$ and $\textbf{M}_{O}^{R}$ are the final left- and right-hand object predictions.
The proposed ORD strategy implicates the physical meaning of the three kinds of objects, especially the meaning of the two-hand objects, i.e., objects that interact with both hands simultaneously. Furthermore, in this way, the network can be trained only to predict the objects that interact with the left-/right-hand rather than focusing on classifying the category of the objects, which achieves decoupling the objects and decreasing the confusion and ambiguity of the algorithm.
III-F Training and inference
The overall ORMNet is trained by the weighted cross-entropy loss function between predictions and GTs of hands, left-hand objects, right-hand objects, and contact boundaries. The overall loss function $\mathscr{L}$ can be computed by,
$\mathscr{L}=\alpha(l_{ce}(\textbf{M}_{O}^{L},\hat{\textbf{G}}_{lo})+l_{ce}(\textbf{M}_{O}^{R},\hat{\textbf{G}}_{ro}),)+$ (14)
$\gamma l_{ce}(\textbf{M}_{H},{\textbf{G}}_{h})+\lambda l_{ce}(\textbf{M}_{CB},\textbf{G}_{cb}),$
where the $l_{ce}(x,y)$ means the cross entropy loss between $x$ and $y$, the $\textbf{M}_{CB}$ and $\textbf{G}_{cb}$ are the prediction and GT of the contact boundary, respectively. The GT of the contact boundary is generated by the GT of the hands and interacting objects automatically, which is the overlapping region of the dilated hand and object masks [4].
For inference, as stated in Sec. III-E, the network predicts the masks of hands, left-hand objects, and right-hand objects, and then the mask of the two-hand objects can be obtained by the proposed ORD strategy.
TABLE I: Comparison results on the EgoHOS in-domain test set measured by IoU/Acc and mIoU/mAcc. The ``"`$\prime$"` ` ′ " means the method with data augmentation, the ``"``"`$\diamondsuit$"` ` ♢ " means the method introduced contact boundary, and the ``"``"`$\flat$"` ` ♭ " means the method with data augmentation and contact boundary.
Method Type Backbone Left hand Right hand Left-hand objects Right-hand objects Two-hand objects Overall
IoU(%)||\,|\,|Acc(%) IoU(%)||\,|\,|Acc(%) IoU(%)||\,|\,|Acc(%) IoU(%)||\,|\,|Acc(%) IoU(%)||\,|\,|Acc(%) mIoU(%)||\,|\,|mAcc(%)
Segformer[13] parallel MiT[13] 62.49||\,|\,|75.47 64.77||\,|\,|78.13 4.03||\,|\,|4.57 3.01||\,|\,|3.17 5.13||\,|\,|5.57 27.89||\,|\,|33.38
SCTNet[54] parallel SCTNet[54] 81.94||\,|\,|90.25 82.12||\,|\,|89.92 17.77||\,|\,|24.49 16.60||\,|\,|20.79 21.74||\,|\,|29.08 44.03||\,|\,|50.91
[…]
Segformer[13] parallel Swin-B 65.36||\,|\,|78.20 66.26||\,|\,|79.21 11.63||\,|\,|15.45 8.07||\,|\,|9.24 12.48||\,|\,|15.97 32.76||\,|\,|39.61
Para.[4] parallel Swin-B 69.08||\,|\,|75.57 73.50||\,|\,|75.93 48.67||\,|\,|39.50 36.21||\,|\,|39.33 37.46||\,|\,|42.58 52.98||\,|\,|54.58
Segmenter[37] parallel Swin-B 82.20||\,|\,|89.87 83.28||\,|\,|91.92 46.22||\,|\,|62.69 34.79||\,|\,|45.59 51.10||\,|\,|62.78 59.52||\,|\,|70.57
UperNet[16] parallel Swin-B 89.88||\,|\,|89.86 91.39||\,|\,|91.32 36.22||\,|\,|37.24 40.55||\,|\,|42.26 45.54||\,|\,|49.27 60.71||\,|\,|61.99
Segmenter[37] parallel ViT[34] 88.47||\,|\,|95.66 89.29||\,|\,|95.08 49.87||\,|\,|61.14 40.60||\,|\,|57.07 46.96||\,|\,|60.81 63.04||\,|\,|73.95
MaskFormer[38] parallel Swin-B 90.45||\,|\,|95.90 91.95||\,|\,|96.41 43.51||\,|\,|67.08 41.04||\,|\,|52.91 54.65||\,|\,|64.86 64.32||\,|\,|75.43
Mask2Former[39] parallel Swin-B 90.74||\,|\,|96.01 92.25||\,|\,|96.20 44.22||\,|\,|53.97 46.05||\,|\,|58.10 51.13||\,|\,|60.48 64.88||\,|\,|72.95
Seq.[4] sequential Swin-B 73.17||\,|\,|- 80.56||\,|\,|- 54.83||\,|\,|- 38.48||\,|\,|- 39.14||\,|\,|- 57.24||\,|\,|-
[…]
ORMNet parallel Swin-B 92.34 ||\,|\,|96.64 93.64 ||\,|\,|96.81 60.07 ||\,|\,|71.79 56.69 ||\,|\,|68.71 54.73 ||\,|\,|65.85 71.49 ||\,|\,|79.96
IV Experiments
This section first introduces the datasets and metrics we used to validate the EgoHOS performance. Second, we compare the proposed ORMNet with state-of-the-art methods in the EgoHOS task. Then, we ablate the import design strategy and elements of ORMNet.
IV-A Datasets and Metrics
EgoHOS. The EgoHOS [4] dataset consists of 11,743 egocentric images containing per-pixel segmentation labels of hands and interacting objects, gathered from Ego4D[55], EPICKITCHEN[56], THU-READ[57], and their own collected egocentric videos. Among them, 8,993 images are used for training, 1,124 images are for validation, 1,126 images are used for in-domain testing, and 500 images are used for out-of-domain testing.
mini-HOI4D. HOI4D[58] is a large-scale egocentric image and video dataset with rich annotations for category-level hand-object interaction. Frame-wise annotations for action segmentation, motion segmentation, panoptic segmentation, 3D hand pose, and category-level object pose are also provided, together with reconstructed object meshes and scene point cloud. In order to test the generalization of the proposed ORMNet, this paper further tests the proposed model with other comparative methods on the HOI4D dataset. Since the HOI4D dataset includes the masks of the hand and all objects that have interacted with the hand in a video, we first extract the masks of the objects currently interacting with the hand and delete other objects’ masks. Then, we map the label of the interacting objects to the left-hand objects, right-hand objects, and two-hand objects. Finally, we sampled a total of 1,095 egocentric images, forming a new dataset mini-HOI4D, to test the generalization of the algorithms on the EgoHOS task.
Following most related works, this paper uses the most common and widely used metrics in the segmentation task to evaluate the performance, i.e., we use Intersection-over-Union (IoU) and pixel Accuracy (Acc) for each category. Mean IoU (mIoU) and mean accuracy (mAcc) are also employed to indicate the overall performance.
IV-B Implementation Details
The experiments were conducted on 4 NVIDIA RTX 6000 Ada GPUs. The batch size was set to 12. Data preparation included cropping the image to (448, 448) and normalization with the mean of [106.011, 95.400, 87.429], and the standard deviation was set to [64.357, 60.889, 61.419]. The patch size of the Swin transformer encoder was set to 4, and the window size was set to 12. The AdamW optimizer was used in the experiment, the learning rate was set to 1e-5, and the weight decay was set to 0.01.
IV-C Quantitative Results
1) Comparison results on EgoHOS in-domain test set. Table I presents the comparison results of our proposed ORMNet against state-of-the-art methods on the EgoHOS in-domain test set. In addition to the parallel (Para.-series) and sequential methods (Seq.-series) proposed in Zhang’s work[4] introduced in detail in Sec. II-B, we also select some of the latest transformer-based models with different backbones as parallel methods to compare the effectiveness of the model, e.g., Segformer[13], Segmenter[37], MaskFormer[38], and Mask2former[39], and SCTNet[54]. All these methods are trained on the EgoHOS train set and tested on the in-domain test set using the same parameter setting as the proposed ORMNet. In general, the proposed ORMNet far outperforms all previous methods in most categories on both metrics. Specifically, compared with the previous best parallel method in the $8^{th}$ row, Mask2Former, our method has demonstrated pronounced advancements across various categories and exhibited particularly notable enhancements in object segmentation. For example, our method showcases substantial enhancements of 15.85%, 10.64%, and 3.6% on left-, right-, and two-hand objects, respectively. Furthermore, compared with the results of the previous best sequential methods $\text{Seq}^{\flat}$, our ORMNet improves IoU by more than 4% on both the left and right hands. The most substantial enhancement is observed in the right-hand objects, surpassing 12%. Finally, compared with the previous best method, our proposed method attains a commendable advancement of 4.33% and 4.55% in mIoU and mAcc, respectively. This is because we propose an ORD strategy to decouple the relationship between object categories, which reduces the burden of network learning. Furthermore, we propose the HOR module to strengthen the correlation between hand features and object features, significantly improving the segmentation accuracy of hands and objects.
TABLE II: Comparison results on the egohos out-of-domain test set.
Method Type Backbone Left hand Right hand Left-hand objects Right-hand objects Two-hand objects Overall
IoU(%)||\,|\,|Acc(%) IoU(%)||\,|\,|Acc(%) IoU(%)||\,|\,|Acc(%) IoU(%)||\,|\,|Acc(%) IoU(%)||\,|\,|Acc(%) mIoU(%)||\,|\,|mAcc(%)
Segformer[13] parallel MiT[13] 71.97||\,|\,|87.33 71.44||\,|\,|81.75 7.60||\,|\,|8.34 5.00||\,|\,|5.65 4.91||\,|\,|5.95 32.18||\,|\,|37.80
Segformer[13] parallel Swin-B 74.01||\,|\,|86.80 70.84||\,|\,|79.44 15.61||\,|\,|21.77 7.32||\,|\,|9.56 8.29||\,|\,|11.67 35.21||\,|\,|41.85
SCTNet[54] parallel SCTNet[54] 87.12||\,|\,|94.62 86.29||\,|\,|90.92 31.18||\,|\,|49.14 19.70||\,|\,|27.47 13.32||\,|\,|17.12 47.52||\,|\,|55.85
UperNet[16] parallel Swin-B 93.17||\,|\,|96.89 93.96||\,|\,|96.00 42.53||\,|\,|64.83 28.88||\,|\,|54.59 24.35||\,|\,|27.83 56.58||\,|\,|68.03
Segmenter[37] parallel Swin-B 83.13||\,|\,|91.22 84.85||\,|\,|92.56 57.97||\,|\,|72.44 38.59||\,|\,|52.75 44.98||\,|\,|52.88 61.90||\,|\,|72.37
Maskformer[38] parallel Swin-B 92.69||\,|\,|95.58 94.02||\,|\,|96.10 51.81||\,|\,|70.53 39.84||\,|\,|60.49 39.43||\,|\,|46.52 63.56||\,|\,|73.84
Mask2former[39] parallel Swin-B 91.46||\,|\,|97.05 93.04||\,|\,|96.38 53.41||\,|\,|64.39 44.90||\,|\,|64.18 35.61||\,|\,|39.78 63.68||\,|\,|72.36
Segmenter[37] parallel ViT[34] 89.40||\,|\,|95.02 90.58||\,|\,|94.86 52.73||\,|\,|75.75 43.88||\,|\,|56.34 42.33||\,|\,|51.08 63.78||\,|\,|74.61
[…]
ORMNet parallel Swin-B 94.47 ||\,|\,|97.09 94.41 ||\,|\,|96.69 51.56 ||\,|\,|72.30 36.80 ||\,|\,|60.90 41.84 ||\,|\,|46.28 63.82 ||\,|\,|74.65
TABLE III: Comparison results on the mini-HOI4D test set. Note: The left-hand objects are not considered to compare due to few samples.
Method Type Backbone Left hand Right hand Right-hand objects Two-hand objects Overall
IoU(%)||\,|\,|Acc(%) IoU(%)||\,|\,|Acc(%) IoU(%)||\,|\,|Acc(%) IoU(%)||\,|\,|Acc(%) mIoU(%)||\,|\,|mAcc(%)
Segformer[13] parallel Swin-B 27.73||\,|\,|87.97 53.27||\,|\,|66.91 8.43||\,|\,|9.72 9.30||\,|\,|13.40 24.68||\,|\,|44.50
Segformer[13] parallel MiT[13] 30.16||\,|\,|92.13 56.44||\,|\,|72.42 5.17||\,|\,|5.52 12.02||\,|\,|13.41 25.95||\,|\,|45.87
SCTNet[54] parallel SCTNet[54] 35.83||\,|\,|95.25 66.29||\,|\,|71.27 17.72||\,|\,|22.68 20.98||\,|\,|29.98 35.21||\,|\,|54.80
Segmenter[37] parallel Swin-B 70.31||\,|\,|89.02 73.78||\,|\,|89.23 21.82||\,|\,|43.75 44.98||\,|\,|56.97 52.72||\,|\,|69.74
UperNet[16] parallel Swin-B 54.82||\,|\,|97.71 84.43||\,|\,|86.04 20.34||\,|\,|25.77 29.34||\,|\,|36.58 47.23||\,|\,|61.53
MaskFormer[38] parallel Swin-B 58.50||\,|\,|96.63 83.66||\,|\,|87.83 35.28||\,|\,|44.81 56.91||\,|\,|73.47 58.59||\,|\,|75.69
Segmenter[37] parallel ViT[34] 74.70||\,|\,|94.03 85.58||\,|\,|92.34 22.38||\,|\,|44.42 58.67||\,|\,|73.99 60.33||\,|\,|76.20
Mask2Former[39] parallel Swin-B 70.13||\,|\,|97.48 88.57||\,|\,|89.38 32.37||\,|\,|45.22 55.72||\,|\,|74.17 61.70||\,|\,|76.56
[…]
ORMNet parallel Swin-B 70.39 ||\,|\,|97.79 89.76 ||\,|\,|93.09 27.56 ||\,|\,|49.35 60.08 ||\,|\,|68.01 61.95 ||\,|\,|77.06
2) Comparison results on EgoHOS out-of-domain test set. In order to test the generalization performance of different models, the EgoHOS dataset provides 500 out-of-domain images, which are captured from YouTube and annotated manually. Table II illustrates the comparison results of our method with the previous methods on the EgoHOS out-of-domain test set measured by IoU and Acc. Please note that some of the results in the table are from direct inference using the officially released best model (if available), and other results are from direct inference on the best model trained on our own. It can be observed that the results of our proposed ORMNet on the two metrics of most categories have obtained competitive generalization performance. Among them, the IoU of the left and right hand exceed 94%, and the Acc. of these two categories exceed 96%. The segmentation results on various object categories also achieve balanced performance. Finally, our ORMNet achieves state-of-the-art performance in overall mIoU and mAcc metrics and demonstrates excellent generalization performance on the out-of-domain test set.
3) Comparison results on mini-HOI4D dataset. In addition to the evaluation on the EgoHOS out-of-domain test set, we also assess the model’s generalization capabilities using a completely different mini-HOI4D dataset. It should be noted that due to the limited samples of left-hand objects in the original HOI4D[58] dataset, the generated mini-HOI4D dataset does not include samples of left-hand objects. Consequently, no testing is performed on left-hand objects.
The comparison results on the mini-HOI4D data set are shown in Table III. Similarly to the EgoHOS out-of-domain test set, the results on the mini-HOI4D dataset are directly inferred by utilizing trained and saved checkpoints without further training procedures. Overall, our proposed ORMNet outperforms all the previous parallel and sequential methods concerning mIoU and mAcc. Our method excels notably in hand segmentation, i.e., 1.57% and 3.71% improvements on IoU and Acc for the right hand compared with the $8^{th}$ row. Our method also competes well in interacting object segmentation, yielding commendable outcomes. Therefore, it can be deduced from the results of the EgoHOS out-of-domain and mini-HOI4D test sets that our proposed ORMNet exhibits superior generalization capabilities in contrast to the comparative methods.
[…]
TABLE IV: Ablation results on the EgoHos in-domain test set under the same setting measured by IoU and mIoU. Lh: left hand, rh: right hand, lo: left-hand object, ro: right-hand object, to: two-hand object.
Method Lh (%) Rh (%) Lo (%) Ro (%) To (%) mIoU (%)
Basic 90.47 91.02 44.11 45.57 51.08 64.45
Basic+ORD 93.14 93.73 60.86 54.58 49.69 70.40
Basic+HOR 92.76 93.36 44.15 53.68 51.85 67.16
Basic+ORD+HOR 92.34 93.64 60.07 56.69 54.73 71.49
TABLE V: The experimental results using different hyper-parameters. $\alpha$, $\gamma$, and $\lambda$ are the weights of objects, hands, and CB losses. Lh: left hand, rh: right hand, lo: left-hand object, ro: right-hand object, to: two-hand object.
$\alpha$ $\gamma$ $\lambda$     Lh (%) Rh (%) Lo (%) Ro (%) To (%)     mIoU (%)
1.0 1.0 1.0     90.66 91.80 57.38 56.20 52.12     69.63
1.0 0.5 0.5     89.74 90.56 57.8 55.46 51.08     68.93
0.5 0.5 1.0     90.22 91.10 57.78 55.59 50.81     69.10
0.5 1.0 0.5     92.34 93.64 60.07 56.69 54.73     71.49
4) Ablation study. In order to verify the effectiveness of the proposed method alongside its corresponding sub-module and strategy, we conduct the ablation study in this section. The experimental results are shown in Table IV. The basic model uses Swin-B as the encoder and uses three decoder branches to predict hands, contact boundaries, and objects. This basic model is a naive solution for the EgoHOS task and does not model any relationships between hands and interacting objects, so its results are relatively poor. The $2^{nd}$ and $3^{rd}$ rows of the table show the results after adding the ORD strategy and HOR module independently, which proves that adding them to the basic model is beneficial. The last row is the result after integrating the ORD strategy and the HOR module. By explicitly modeling both the hand-object relationship and the object-object relationship, the segmentation accuracy achieves its peak performance. Generally speaking, although the effect improvement brought by the ORD strategy is higher than that of the HOR module, using these two modules and strategies simultaneously can further improve the segmentation performance, which verifies the effectiveness of our proposed method and sub-methods.
Figure 5: Visualization results of the ORMNet compared with the sequential method on the EgoHOS out-of-domain test set (left) and mini-HOI4D dataset (right). Red: left hand, dark blue: right hand, pink: left-hand objects, light blue: right-hand objects, and green: two-hand objects.
5) Hyper-parameter experiments. While identifying the optimal hyper-parameters is not the primary focus of this paper, this section still conducts experiments on the configurations of different hyper-parameters. The experimental results are shown in Table V. We assign the value of $\alpha$, $\gamma$, and $\lambda$ based on the relative significance of the interacting objects, hands, and contact boundaries. In the $1^{st}$ row, all categories are considered equal importance, while in subsequent experiments (i.e., $2^{nd},3^{rd}$ and $4^{th}$), we regard the objects, CBs, and hands as the most crucial, respectively. The results lead to the inference that amplifying the importance of hands can yield enhancements in the method’s performance, which can be explained intuitively because the proposed HOR module leverages hands as prior knowledge for feature enhancement.
IV-D Qualitative Results
[…]
2) Results on EgoHOS out-of-domain test set. We directly used the saved checkpoint to conduct experiments on the EgoHOS out-of-domain test set and compared it with the $\text{Seq}^{\flat}$ method. The visualized experimental results are shown in Fig. 5 (left). According to the visualization results, our ORMNet has stronger generalization than the compared sequential model. Specifically, our method can segment and identify left and right hands accurately, as shown in the $3^{rd}$ to the last rows of results. Furthermore, our method can precisely segment and recognize different kinds of objects, ensuring consistent segmentation of interactive objects.
3) Results on mini-HOI4D dataset. The experiments on the mini-HOI4D dataset are also conducted directly using the saved checkpoints. The experimental results are shown in Fig. 5 (right). It can be observed from the experimental results that the contrasting sequential method easily confuses the left and right hands, such as $1^{st}$, $2^{nd}$, $4^{th}$, and $5^{th}$ rows of the results. In contrast, our method exhibits robust generalization capabilities coupled with outstanding segmentation performance. In addition, although some tiny parts are not considered, such as the handle of the bucket in the third row, our method also significantly excels in segmenting two-hand objects.
V Conclusion
This paper introduces an end-to-end high-performance Object-centric Relationship Modeling Network to effectively handle the egocentric hand-object segmentation task. First, the HOR module is proposed to facilitate the modeling of hand-object interactions through hand-guided cross-attention, enhancing the network’s adaptability to learn precise representations. Second, this paper designs the ORD strategy to decouple the two-hand objects, thereby reducing the network’s confusion about object segmentation. Experimental results demonstrate that the proposed ORMNet exhibits SOTA segmentation efficacy and robust generalization capabilities across varied datasets.
VI Acknowledgements
The research work was conducted in the JC STEM Lab of Machine Learning and Computer Vision funded by The Hong Kong Jockey Club Charities Trust.

### v2

@@
 Abstract
-Egocentric hand-object segmentation (EgoHOS) is a brand-new task aiming at segmenting the hands and interacting objects in the egocentric image.
-Although significant advancements have been achieved by current methods, establishing an end-to-end model with high accuracy remains an unresolved challenge.
-Moreover, existing methods lack explicit modeling of the relationships between hands and objects as well as objects and objects, thereby disregarding critical information on hand-object interaction and introducing confusion into algorithms, ultimately leading to a reduction in segmentation performance.
-To address the limitations of existing methods, this paper proposes a novel end-to-end Object-centric Relationship Modeling Network (ORMNet) for EgoHOS.
-Specifically, based on a single-encoder and multi-decoder framework, we design the Hand-Object Relation (HOR) module to leverage hand-guided attention to capture the correlation between hands and objects and facilitate their representations.
-Moreover, based on the observed interrelationships between diverse categories of objects, we introduce the Object Relation Decoupling (ORD) strategy.
-This strategy allows the decoupling of the two-hand object during training, thereby alleviating the ambiguity of the network.
-Experimental results on three datasets show that the proposed ORMNet has notably exceptional segmentation performance with robust generalization capabilities.
-Index Terms:
-Egocentric view, hand-object interaction segmentation, attention, relationship modeling.
-I Introduction
-The advancements in edge computing and computer vision have engendered a growing interest in Head-Mounted Devices (HMD) such as the Apple Vision Pro[1] and Microsoft Hololens 2[2].
-Consequently, there is a rising trend in capturing abundant first-person view (FPV) or egocentric images and videos.
-Compared with data captured from a third-person view (TPV) or an exocentric perspective[3], egocentric data displays the actions and intentions of the HMD wearer more clearly and intuitively, which is more conducive to analyzing human behavior and understanding how humans interact with the environment.
-Notably, questions concerning the activities of individuals, such as “What is this person doing?” can be more comprehensively addressed through the parsing of egocentric visual data.
+Egocentric hand-object segmentation (EgoHOS) is a promising new task aiming at segmenting hands and interacting objects in egocentric images.
+Although EgoHOS has the potential to enable various applications, current methods struggle to achieve both high performance and end-to-end optimization simultaneously.
+Moreover, existing approaches fail to fully leverage hand cues to assist the interacting-object segmentation and overlook the coupled relationships between diverse interacting-object categories, resulting in performance deficiencies.
+To address these limitations, this paper proposes a novel Object-centric Relationship Modeling Network (ORMNet) to fulfill end-to-end and effective EgoHOS by modeling relationships between hands and objects as well as objects and objects.
+Specifically, a Hand-Object Relation (HOR) module is introduced to capture the correlation between hands and objects, which uses hand features to guide the network to extract more distinguishing interacting-object features.
+Besides, we find the coupling relations between diverse interacting-object categories and design the Object Relation Decoupling (ORD) strategy to disentangle them, emphasizing learning of the interaction between hands and objects and reducing the confusion of interacting-object classification.
+In-domain experiments show that ORMNet has notably exceptional segmentation performance compared with state-of-the-art methods, while out-of-domain experiments further exhibit its robust generalization capability.
+The project is available at https://github.com/yuggiehk/ORMNet/ .
+keywords:
+Egocentric perception , hand-object segmentation , cross-attention , relationship modeling
+journal: Pattern Recognition
+[…]
+[label1]organization=Department of Electrical and Electronic Engineering, The Hong Kong Polytechnic University, addressline=Hung Hum, city=Kowloon, country=Hong Kong SAR
+1 Introduction
+The advancements in edge computing and computer vision [1, 2, 3] have engendered a growing interest in Head-Mounted Devices (HMD) such as the Apple Vision Pro [4] and Microsoft Hololens 2 [5].
+Consequently, there is a rising trend in capturing abundant first-person view (FPV) or egocentric images and videos [6, 7].
+Compared with data captured from a third-person view (TPV) or an exocentric perspective [8], egocentric data displays more detailed, intuitive, and clear actions and intentions of the HMD wearer, which is more conducive to analyzing human behavior and understanding how people interact with the environment.
+Notably, questions concerning the activities of individuals, such as “What is this person doing?” can be more comprehensively addressed in detail through the parsing of egocentric visual data.
 It is obvious that the hands of HMD wearers and the objects they are interacting with serve as pivotal cues in this analysis.
-Therefore, to delve deeply into human behavior encoded within egocentric data at a fine-grained level, Zhang et al.[4] first proposed the Egocentric Hand-Object Segmentation (EgoHOS) task, which aims to perform pixel-level segmentation of the hand and the objects that interact with the hand in egocentric images.
-This novel segmentation task enables a precise response to the query, “Where are the left and right hands, and what objects are they interacting with?” which can function as an essential algorithm serving as the foundation for various subsequent applications, such as augmented reality/virtual reality (AR/VR)[5, 6], medical assistant[7, 8], robotics[9], embodied AI[10, 11], and metaverse[12].
-Figure 1: Structural differences between the previous (a) parallel method, (b) sequential method, and (c) our approach.
-Compared with previous parallel and sequential methods, the proposed ORMNet can achieve higher segmentation performance with higher efficiency by integrating relationship modeling between “hands and objects (H-O)” as well as “objects and objects (O-O).” The $N,N_{h},$ and $N_{o}$ mean the number of overall categories, the number of hand categories, and the number of object categories, and $N=N_{h}+N_{o}$.
+Therefore, to delve deeply into human behavior encoded within egocentric data at a fine-grained level, Zhang et al.
+[9] first proposed a new task, i.e., Egocentric Hand-Object Segmentation (EgoHOS), which aims to perform pixel-level segmentation of hands and objects that are interacting with hands in egocentric images.
+This novel segmentation task enables a precise response to the query, “Where are the left and right hands, and what objects are they interacting with?” which can function as an essential algorithm serving as the foundation for various subsequent applications, such as augmented reality/virtual reality (AR/VR) [10, 11], medical assistive systems [12, 13], and embodied AI [14, 15].
+Figure 1: (a) parallel methods [9, 16, 17, 18, 19, 40], (b) sequential methods [9], and (c) our approach.
+Comparisons of segmentation performance (d) on the EgoHOS in-domain and out-of-domain test sets show that our ORMNet can achieve significant segmentation and generalization performance with end-to-end optimization by integrating the HOR module and ORD strategy.
 The previous approaches for EgoHOS can be categorized into parallel and sequential methods.
-Specifically, parallel methods[4, 13, 14, 15, 16] are straightforward, which employ an end-to-end encoder-decoder framework to extract features and predict the mask of all categories simultaneously, as shown in Fig.
-1(a).
-In contrast to parallel methods, sequential approaches[4, 17] employ a step-by-step prediction scheme for the hand and interacting objects, i.e., the input of the latter step to predict the objects’ masks necessitates the previous step’s output for predicting the hand, as depicted in Fig.
-1(b).
-Although existing methods have demonstrated a satisfied EgoHOS capability, they are accompanied by many drawbacks, as shown in the table in Fig.
+Specifically, some unified segmentation methods [9, 16, 17, 18, 19] can be directly used as parallel methods in this new task, which adopt an end-to-end framework to first extract image features and then simultaneously predict all categories of masks, as shown in Fig.
+1 (a).
+However, due to the newly-emerging EgoHOS task, applying such generic segmentation methodologies directly to this task lacks targeted design of relation modeling, resulting in suboptimal effectiveness.
+In contrast to parallel methods, sequential approaches [9] are specially designed to form a step-by-step prediction scheme for hands and interacting objects, i.e., the input of the latter step to predict the objects’ masks necessitates the previous step’s output for predicting hands, as depicted in Fig.
+1 (b).
+Although the effectiveness of sequential methods surpasses that of parallel methods, they are unable to achieve end-to-end optimization, which reduces the efficiency of the network.
+Therefore, achieving both high performance and end-to-end optimization in this new EgoHOS task remains an unresolved issue.
+Furthermore, regardless of variations in interactions and environments, it is easier to recognize hands in egocentric images by identifying the shape and texture [20, 7].
+Conversely, the segmentation of interacting objects presents significant challenges due to the reliance on discerning interaction with hands rather than visual features.
+Therefore, the dominant role of hands in the EgoHOS task underscores the benefit of incorporating hands to assist in predicting interacting objects.
+However, as mentioned above, parallel methods [9, 21, 22] treat hands and interacting objects as completely independent categories for prediction, while sequential methods [9] use predicted hand masks as input for interacting-object segmentation.
+Both approaches fail to subtly and explicitly model the dominance of hands in predicting interacting objects, thereby leading to performance deficiencies.
+In addition, the EgoHOS task requires the prediction of three types of interacting objects determined by the hand(s) they interact with, namely, left-hand objects, right-hand objects, and two-hand objects, as described in detail in Sec.
+3.1.
+We observe that the two-hand object encompasses attributes of both the left- or right-hand object because they are interacting with both left- and right-hand simultaneously.
+However, most existing parallel and sequential approaches treat different types of interacting objects as separate, which requires the network to learn “whether the object is interacting with both hands at the same time,” introducing additional confusion into the training process and resulting in inaccurate results.
+To address the shortcomings of the existing methods above, this paper proposes a novel Object-centric Relationship Modeling Network (ORMNet) to model the relationships between “hands and objects” as well as “objects and objects”, which can achieve end-to-end optimization with significant performance, as shown in Fig.
+1 (c).
+On the one hand, we propose the Hand-Object Relation (HOR) module to establish the correlation between the features of the hand and interacting objects.
+The core underlying concept involves leveraging hand features as prior knowledge and employing hand-guided attention to extract more representative interacting-object features.
+On the other hand, based on the observations about interacting objects, we introduce a novel Object Relation Decoupling (ORD) strategy to disentangle the interrelationships among diverse interacting-object categories.
+This strategy restricts the prediction of the left- and right-hand objects exclusively during training, and another type of object (two-hand objects) is generated through an ingenious and pure mathematical intersection operation during inference.
+This strategy can help the algorithm emphasize the interaction and reduce the confusion of the algorithm about the interacting-object classification during learning.
+Experimental results on in-domain and out-of-domain test sets demonstrate the state-of-the-art performance of ORMNet along with robust generalization capabilities, as shown in Fig.
+1 (d).
+The contributions of this paper are as follows:
   • 1.
-First, similar to conventional encoder-decoder semantic segmentation networks[18, 19], the intuitive parallel approaches[4] regard hands and objects as independent categories and predict segmentation results for them in an end-to-end manner.
-While this type of technique demonstrates high efficiency, it fails to capture the relationship between “hands and objects” or among different objects, consequently leading to performance deficiencies.
-Second, in contrast, the sequential methods[4] specially built for the EgoHOS task demonstrate enhanced performance by incorporating the relationships between hands and interacting objects.
-However, step-wise training and inference elongate the overall time consumption, thereby reducing the method’s efficiency and constraining its feasibility for real-time applications.
-Third, the prediction of diverse interacting objects is required in the EgoHOS task, i.e., left-hand objects (objects interact with the left hand), right-hand objects (objects interact with the right hand), and two-hand objects (objects interact with both hands at the same time).
-This paper observes the coupling relationship between these three categories of objects, i.e., the two-hand object can be regarded as both a left-hand object and a right-hand object.
-Therefore, it is easy to predict a two-hand object as the left-hand object when the left interaction is more obvious than the right interaction, and the same applies to the right-hand object.
-Nevertheless, prevailing methods overlook the ambiguity of such relationships, which introduces confusion during network training.
-To improve efficiency and effectiveness simultaneously, this paper proposes a novel end-to-end Object-centric Relationship Modeling Network (ORMNet) to model the relationship between “hands and objects” as well as “objects and objects,” as shown in Fig.
-1(c).
-On the one hand, we propose the Hand-Object Relation (HOR) module to establish the correlation between the features of the hand and interacting objects.
-The core underlying concept involves leveraging hand features as prior knowledge and employing hand-guided attention to extract more representative object features.
-On the other hand, we introduce a new Object Relation Decoupling (ORD) strategy to model the relationship among diverse object categories.
-Drawing from the insights gained from analyzing two-hand objects, this strategy restricts the prediction to left- and right-hand objects exclusively during the training phase.
-Then, two-hand objects are generated through a simple yet efficient intersection operation during inference.
-This strategy can reduce the confusion of the algorithm about the object categories during learning, enabling the network to concentrate more precisely on the interacting objects.
-The contributions of this paper are as follows:
-  • This paper introduces a novel end-to-end ORMNet for the EgoHOS task that achieves state-of-the-art segmentation performance while being highly efficient.
-  • To take full advantage of the prominence of hand features, the proposed HOR module models the relationship between hands and objects, helping the network extract more comprehensive hand and object features.
-  • Based on the relation of various object categories, the simple yet powerful ORD strategy is proposed to decouple the two-hand objects during training, reducing the ambiguity of the network and further separately highlighting the objects contacted with left or right hands.
-  • The experimental results on two datasets demonstrate that the proposed ORMNet achieves exceptional performance and manifests robust generalization capabilities.
-II Related Works
-II-A Image Segmentation
-Image segmentation is a basic task of computer vision, which aims to divide the image into several non-overlapping regions based on certain principles, such as semantic information.
-Semantic segmentation[20, 21, 22], instance segmentation[23, 24, 25], and panoptic segmentation[26, 27, 28] are three commonly studied segmentation sub-tasks.
-Image segmentation methods can be divided into traditional methods, convolution neural network (CNN)-based methods, and transformer-based methods in chronological order.
-Although traditional methods can perform image segmentation with notable interpretability, further refinement is needed to enhance the quality of the segmentation outcomes.
-The fully connected neural network (FCN)[19] proposed by Jonathan et al.
-is the cornerstone of CNN segmentation methods.
-After that, there have been numerous advancements[29] in CNN-based image segmentation methods, such as U-Net[18], Deeplab-series[30, 31], SegNet[32], PSPNet[33], etc.
-The development of transformer-based methods for image segmentation is a relatively recent but rapidly evolving field.
-Dosovitskiy et al.
-introduced the Vision Transformer (ViT)[34] in 2020, proving that the transformer could be applied directly to images by treating the image as a sequence of patches.
-Following the success of ViT, several transformer-based methods[35] specially designed for image segmentation have been proposed.
-Zheng et al.
-developed the SETR[36] to adopt ViT as a backbone to extract features to eliminate the reliance on FCN, achieving impressive performance.
-Similarly, Strudel et al.
-proposed a pure end-to-end Transformer approach - Segmenter[37], which utilized the ViT encoder for capturing global context and a simple point-wise linear decoder to generate precise segmentation masks.
-Meanwhile, Xie et al.
-presented the Segformer[13] to integrate hierarchical Transformer encoder with lightweight multi-layer perceptron (MLP) decoders, avoiding complex designs in segmentation methods.
-Unlike previous work, Cheng et al.
-proposed the MaskFormer[38], which abandoned the per-pixel classification formulation and instead predicted a series of binary masks associated with class labels.
-Following that, Cheng et al.
-introduced the Mask2Former[39], which incorporated masked attention to perform universal (i.e., instance, semantic, and panoptic) segmentation, achieving notable performance.
-Then, Jain et al.
-proposed OneFormer[14] to perform universal segmentation tasks with transformers and task-guided queries, which achieved state-of-the-art performance in various segmentation benchmarks.
-Recently, Kirillov et al.
-proposed a new impressive work called Segment Anything (SAM)[40].
-They built a foundation model for interactive segmentation by training with over 1 billion masks on 11M images.
-Our paper utilizes the Swin transformer[41] to extract the features of the input egocentric image and to up-sample the features to the input size[15].
-Although using more recent methods may improve performance, this is not the main focus of this paper.
-II-B Egocentric Hand-object Segmentation
-Zhang et al.[4] recently introduced a novel task known as Egocentric Hand-object Segmentation (EgoHOS), which combines image segmentation and Human-Object Interaction detection (HOI-det).
-This task redefines the focus of HOI-det from the entire human body to specific, fine-grained parts, i.e., human hands.
-Its objective is to segment the left hand, right hand, and the objects that interact with the left and right hands in input RGB egocentric images.
-Concurrently, Zhang et al.
-[4] introduced the innovative concept of the contact boundary (CB), representing the region where the hands and the objects engage in interaction.
-This concept explicitly promotes the network focus on the interaction between hands and objects.
-Furthermore, based on the proposed task and concept, the authors also proposed different types of methods in their work, including parallel methods and sequential methods.
-Parallel method.
-A more direct and naive method is to use an encoder and a decoder to segment and predict hands and objects and use different indexes to imply different categories[4, 13, 14, 15, 16], as shown in Fig.
-1(a).
-Although this parallel method completes the prediction of this task simply and directly, its segmentation accuracy is relatively low and still needs to be improved.
-Sequential method.
-Since parallel methods lacked targeted design in learning hand-object interaction, Zhang et al.
-proposed a sequential method[4], as shown in Fig.
-1(b).
-Taking the egocentric image as input, this step-by-step sequential method uses an encoder-decoder first to segment the left and right hands.
-After that, the segmented left- and right-hand masks are concatenated to the input image to be used to perform the mask prediction of contact boundary and interacting objects.
-Although the sequential method achieved higher segmentation performance than the parallel method, it lacks efficiency due to the step-by-step process.
-In addition, both parallel and sequential methods inadequately explore the relationship between hands and objects and ignore the relationship between objects and objects.
-Unlike the previous parallel and sequential methods, the proposed ORMNet can achieve end-to-end training and inference with high accuracy due to the relationship modeling between “hands and objects” as well as “objects and objects.”
-II-C Attention Mechanism in deep learning
-Models using the attention mechanism have exploded in recent years, confirming the effectiveness of the attention mechanism for NLP[42, 43] and computer vision[44, 45] tasks.
-The primary objective of attention in vision is to mimic the selective function of the human visual cognitive system towards input information, i.e., concentrating on the primary and crucial features of the input visual data.
-The majority of attention modules utilize distinct methodologies to generate an attention map for feature weighting, consequently adjusting the significance of individual components within the features.
-In the early stage, most attention modules are integrated into the CNN-based models.
-For example, Hu et al.
-proposed the SENet[46] equipped with channel-wise attention, which used global average pooling to generate the attention weight among channel axial.
-After that, Woo et al.
-proposed CAM [47] to improve channel attention by aggregating the descriptors from max-pooling and average-pooling.
-There are many other methods intended to improve channel-wise attention, e.g., SOCA[48], HA[49], etc.
-Unlike channel attention, spatial attention is proposed to generate weights for spatial patches of feature maps.
-For example, Woo et al.
-proposed SAM [47] to use a convolutional layer to obtain the attention weights.
-The development of self-attention and multi-head self-attention (MHSA) mechanisms in transformer-based models has also been widely used in various methods in recent years.
-Some researchers also make an effort to extend self-attention to cross-attention[50, 51], which allows the model to capture relationships between different modalities of data[52, 53].
-The proposed ORMNet utilizes hand-guided cross-attention to model the relationship between the features of hands and objects, allowing more comprehensive feature extraction.
-Figure 2: Overall diagram of the proposed ORMNet, which consists of an encoder, bottleneck, multiple decoder branches, Hand-Object Relation (HOR) module, and Object Relation Decoupling (ORD) strategy.
-The algorithm is utilized to predict the hands, right-hand objects (RO), left-hand objects (LO), and two-hand objects (TO).
-The ORD strategy processes the initial (Ini.) ground truth (GT) to generate new supervision during data preparation, and it also processes the predictions (Pred.) of diverse categories of interacting objects during inference.
-III Methodology
-This section introduces the proposed method in detail from top to bottom, including task description (in Sec.
-III-A), overall architecture (in Sec.
-III-B), Object Relation Decoupling Strategy (in Sec.
-III-E), Hand-Object Relation module (in Sec.
-III-D), and training and inference (in Sec.
-III-F).
-III-A Task Description
-The EgoHOS dataset [4] includes the label of the hands, $1^{st}$-order interacting objects, and $2^{nd}$-order interacting objects.
-In this paper, we focus on the segmentation of hands and $1^{st}$-order interacting objects, i.e., objects that are directly contacted by hands.
-Given an egocentric image I, the goal of the EgoHOS task is to predict the masks of the following entities: left hand $\textbf{M}_{H}^{L}$, right hand $\textbf{M}_{H}^{R}$, left-hand objects $\textbf{M}_{O}^{L}$ (objects directly interacting with the left hand), right-hand objects $\textbf{M}_{O}^{R}$ (objects directly interacting with the right hand), and two-hand objects $\textbf{M}_{O}^{T}$ (objects interacting with both hands simultaneously).
-III-B Overall Architecture
+    This paper introduces a novel ORMNet for the EgoHOS task, achieving state-of-the-art segmentation performance while fulfilling end-to-end optimization.
+  • 2.
+    To take full advantage of the prominence of hand features, ORMNet incorporates the HOR module to model the relationship between hands and interacting objects, helping the network extract more representative object features.
+  • 3.
+    Based on the observed relation between various interacting-object categories, an innovative ORD strategy is proposed to decouple them, decreasing the concentration on classification and emphasizing interaction learning of the algorithm.
+  • 4.
+    The thorough experimental results demonstrate that the proposed ORMNet achieves exceptional performance and manifests robust generalization capabilities.
+2 Related Work
+This section mainly introduces the related work of this paper.
+Since the EgoHOS task is strongly related to Egocentric Hand-Object Interaction (EgoHOI), we will first introduce EgoHOI (Sec.
+2.0.1) and then refine to EgoHOS (Sec.
+2.0.2).
+2.0.1 Egocentric Hand-Object Interaction
+Egocentric images and videos offer a clear window into how humans physically engage with their surrounding environment and the objects using their hands.
+In recent years, the interpretation and analysis of egocentric visual data have earned increasing attention from the research community, especially in the field of Egocentric hand-object interaction (EgoHOI) [23, 24, 25] due to the release of various large-scale datasets [26, 24, 27, 28, 29].
+Aiming to solve the EgoHOI task, certain efforts [30] used additional multi-modality information, such as gaze [31] and text [32, 33] cues, to enhance the learnable representations.
+Furthermore, self-supervised diagrams [34] were developed to effectively leverage the underlying visual information.
+Although egocentric visual data is inherently far more limited in quantity compared to the abundant third-person perspective (TPV) data, some works [35, 36] proposed to learn the joint representations and transfer the view-agnostic knowledge from TPV to FPV.
+However, the data requirements are strict, time-consuming, and labor-intensive, which needs videos capturing the same action from different perspectives.
+In addition, EgoHOI can be utilized as a basic algorithm to perform other tasks, such as action recognition [20, 37, 38] and action anticipation [39].
+While many transformer-based approaches have made breakthrough progress on the EgoHOI task, the sole prediction of bounding boxes around the interacting objects lacks the granularity of fine-grained analysis.
+2.0.2 Egocentric Hand-Object Segmentation
+Different from EgoHOI, EgoHOS is a novel problem proposed by Zhang et al.
+[9] to perform fine-grained pixel-level segmentation for hands and interacting objects.
+The final prediction output encompasses a set of categories, including the left hand, right hand, left-hand objects, right-hand objects, and two-hand objects.
+Furthermore, the authors also proposed different types of methods in their work, including parallel methods and sequential methods.
+Parallel methods.
+Some general segmentation methods [9, 16, 17, 18, 19, 40] can be directly used as parallel methods to perform EgoHOS task.
+These methods take egocentric images as input, and directly obtain predictions of all categories after feature extraction and mask generation, as shown in Fig.
+1 (a).
+Although these methods can complete the EgoHOS task in an end-to-end manner, they are not effective due to the lack of task-specific design.
+Sequential methods.
+Zhang et al.
+proposed a series of sequential methods [9] tailored for the EgoHOS task, as shown in Fig.
+1 (b).
+Taking the egocentric image as input, these step-by-step sequential methods first segment the left and right hands.
+After that, the segmented left- and right-hand masks are concatenated to the input image to perform the mask prediction of interacting objects.
+Although sequential methods achieve higher segmentation performance than the parallel method, the step-by-step process is unable to achieve end-to-end optimization, which limits the applications in the real world.
+In addition, both parallel and sequential methods inadequately explore the relationship between hands and interacting objects, which does not emphasize the leading role of the hand [20] in the EgoHOS task, thereby degrading the segmentation performance.
+Finally, parallel and sequential methods regard the different categories of interacting objects as independent and ignore the potential relationships between them, causing unnecessary confusion in classifying the interacting-object categories.
+Unlike the previous parallel and sequential methods, the proposed ORMNet can achieve end-to-end optimization with high accuracy due to the relationship modeling between “hands and objects” as well as “objects and objects,” as shown in Fig.
+1 (c).
+3 Methodology
+This section introduces the proposed ORMNet in detail from top to bottom, including task description (in Sec.
+3.1), overall architecture (in Sec.
+3.2), HOR module (in Sec.
+3.3), ORD strategy (in Sec.
+3.4), and training and inference (in Sec.
+3.5).
+3.1 Task Description
+This section first delineates the detailed description of the newly-emerging EgoHOS task.
+The EgoHOS task includes the segmentation of left- and right-hands, $1^{st}$-order interacting objects, and $2^{nd}$-order interacting objects.
+Following the work proposed by Zhang et al.
+[9], we focus on the segmentation of hands and $1^{st}$-order interacting objects, i.e., objects that are directly contacted by hands.
+Given an egocentric image $\textbf{I}\in\mathbb{R}^{H\times W\times 3}$, the goal is to predict the masks of the following entities: hands $\textbf{M}_{H}$, left-hand objects $\textbf{M}_{O}^{L}$ (objects directly interacting with the left hand), right-hand objects $\textbf{M}_{O}^{R}$ (objects directly interacting with the right hand), and two-hand objects $\textbf{M}_{O}^{T}$ (objects interacting with both hands simultaneously).
+Figure 2: Overall diagram of the proposed end-to-end ORMNet, which incorporates the HOR module and ORD strategy to effectively predict hands, left-hand objects (LO), right-hand objects (RO), and two-hand objects (TO).
+The ORD strategy consists of data preparation (left), training (middle), and inference (right).
+3.2 Overall Architecture
 The Overall diagram of the proposed ORMNet is shown in Fig.
-2.
-The model consists of an encoder, a bottleneck, and a decoder with multiple branches.
-The encoder is used to extract distinguishing features of the input egocentric image, the bottleneck is used to enhance the extracted global features, and different branches of the decoder are used to predict the final segmentation masks for diverse categories.
+2, which consists of data preparation (left), training (middle), and inference (right).
+The detailed information about data preparation and inference will be introduced in Section 3.4.
+The model architecture consists of an encoder, a bottleneck, and a decoder with multiple branches.
+The encoder is used to extract multi-scale distinguishing features of the input egocentric image, the bottleneck is used to enhance the extracted global features, and different branches of the decoder are used to predict the final segmentation masks for diverse categories.
+Specifically, taking an egocentric image $\mathbf{I}$ as input, the first step of ORMNet is to use the image encoder to extract the multi-scale representative features $\textbf{F}=\{\textbf{F}_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times C_{i}}\;|\;i=0,1,\cdots,N\}$ by a series of down-samplings with a factor of 2, where $H_{i}$, $W_{i}$, and $C_{i}$ are the height, width, and channel of the $i^{th}$-stage feature map, respectively.
+The image encoder used in practice is the Swin Transformer encoder [41], and N is set to 3.
+The extraction process of generating multi-scale features in the encoder can be expressed as follows:
+$\mathbf{F}_{i+1}=f_{enc}^{i+1}(\mathbf{F_{i}}\;|\;W_{enc}^{i+1}),\;i\in\left\{0,1,\cdots,N-1\right\},$ (1)
+$\mathbf{F_{0}}=f_{enc}^{0}(\mathbf{I}\;|\;W_{enc}^{0}),$ (2)
+where the $f_{enc}^{i}(\cdot)$ denotes the encoder stage for generating the $i^{th}$-stage feature map, and the $W_{enc}^{i}$ represents corresponding parameters.
+After generating the multi-scale features, the final global feature map $\textbf{F}_{N}$ is sent into the bottleneck to enhance the representation of the encoded feature.
+Following the work [18], the bottleneck is composed of a patch merging layer and two Swin Transformer blocks, which are denoted as follows:
+$\mathbf{F}_{ehc}=f_{bot}(\mathbf{F}_{N}\;|\;W_{bot}),$ (3)
+where the $\mathbf{F}_{ehc}$ represents the enhanced final feature map, and the $f_{bot}(\cdot)$ denotes the bottleneck network with the parameter of $W_{bot}$.
+Taking the enhanced global feature $\mathbf{F}_{ehc}$ and multi-scale features $\textbf{F}^{\prime}=\{\textbf{F}_{i}\;|\;i=0,1,\cdots,N-1\}$ as input, different decoder branches are used to predict the final mask of each category.
 Following the previous work [9], ORMNet also contains the contact boundary (CB) decoder branch to predict the mask of the contact boundary.
-For simplification, we remove the CB decoder branch in Fig.
-2.
-III-C Encoder and Bottleneck
-Taking an egocentric image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ as input, the first step of ORMNet is to use the image encoder to extract the multi-scale features $\textbf{F}_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times C_{i}}$ by a series of down-sampling by a factor of 2, where $H_{i}$, $W_{i}$, and $C_{i}$ are the height, width, and channel of the $i^{th}$ stage feature map, respectively.
-The image encoder used in practice is the Swin Transformer encoder[41].
-The extraction process of generating multi-scale features in the encoder can be expressed as follows:
-$\mathbf{F}_{i+1}=f_{enc}^{i+1}(\mathbf{F_{i}}|W_{enc}^{i+1}),i\in\left\{0,1,2\right\},$ (1)
-$\mathbf{F_{0}}=f_{enc}^{0}(\mathbf{I}|W_{enc}^{0}),$ (2)
-where the $f_{enc}^{i}(\cdot)$ denotes the encoder stage with the down-sampling for generating the $i^{th}$ stage feature map, and the $W_{enc}^{i}$ represents corresponding parameters.
-After generating the multi-scale features, the final feature map with the smallest height and width and the largest channel is sent into the bottleneck to enhance the representation of the encoded feature.
-Following the work[15], the bottleneck is composed of a patch merging layer and two Swin Transformer blocks , which is denoted as follows:
-$\mathbf{F}_{ehc}=f_{bot}(\mathbf{F}_{last}|W_{bot}),$ (3)
-where the $\mathbf{F}_{last}$ and $\mathbf{F}_{ehc}$ represent the final encoded feature map and enhanced final feature map, and the $f_{bot}(\cdot)$ denotes the bottleneck network with the parameter of $W_{bot}$.
-The obtained enhanced memory feature $\mathbf{F}_{ehc}$ and multi-scale features $\{\mathbf{F}_{i}|i=0,1,2\}$ are then passed into separate decoder branches to predict the hands and interacting objects.
-III-D Hand-Object Relation Module
+For simplification, we remove it in Fig.
+2 (middle).
+In each decoder branch, N+1 decoder stages are employed to predict the masks.
+And the output feature of each stage is concatenated with corresponding encoded features by skip-connection [42] as the input to the next stage.
+The final up-sampling layer is used to recover the resolution to the same as the input egocentric image.
+The process of each decoder branch can be denoted as:
+$\textbf{D}_{0}=f_{up}^{0}(\textbf{F}_{ehc}),$ (4)
+$\textbf{D}_{i+1}=f_{up}^{i+1}(f_{cat}(\textbf{D}_{i},\textbf{F}_{N-1-i})),$ (5)
+$\textbf{M}=f_{us}(\textbf{D}_{N}),$ (6)
+where the $f_{up}^{i}(\cdot)$ means the $i^{th}$ decoder stage, which outputs $\textbf{D}_{i}$ as decoded feature map.
+The $f_{cat}(\cdot)$ means the concatenation operation, the $f_{us}(\cdot)$ represents the final up-sampling layer, and the M denotes the predicted mask of this decoder branch.
 Figure 3: Detailed architecture of the proposed HOR module.
-Taking the features of hands and objects as input, this module accomplishes hand-guided attention to model the relationship between the hands and interacting objects, promoting the object feature learning of the network.
-Among the egocentric visual information, the hand information of the HMD wearer is the most prominent and representative.
-Moreover, the interacted objects can only be identified based on whether they are in contact with the hand rather than appearance and texture cues.
-Therefore, using hand features to enhance the extraction and expression of object features is beneficial to the EgoHOS task, which has been ignored in previous work.
-This paper proposes the HOR module to model the correlation between hand features and object features in interaction.
-Specifically, this module uses representative hand features as priors and guidance and utilizes the hand-guided attention mechanism to guide the learning of object features, as shown in Fig.
+Taking the features of hands and interacting objects as input, this module accomplishes hand-guided attention to model the relationship between them, promoting the object feature learning of the network.
+3.3 Hand-Object Relation Module
+Among the egocentric visual information, the hand information of the HMD wearer is the most prominent and representative because we can easily recognize hands by shape and texture [20].
+Moreover, the contacted objects can be accurately identified based on whether they are interacting with hands rather than appearance and texture cues.
+Therefore, using hand features to enhance the extraction and expression of interacting-object features is beneficial to the EgoHOS task.
+While certain sequential methods [9] employ predicted hand masks to segment interacting objects, the performance of this coarse approach cannot effectively excavate the relationships between them.
+This paper proposes the HOR module to model the correlation between hand and object features in interaction, which uses representative hand features as priors and guidance and utilizes the hand-guided attention mechanism to guide the learning of interacting-object features, as shown in Fig.
 3.
-Taking the hand features $\textbf{F}_{hand}\in\mathbb{R}^{\hat{C}\times\hat{H}\times\hat{W}}$ and the object features $\textbf{F}_{obj}\in\mathbb{R}^{\hat{C}\times\hat{H}\times\hat{W}}$ generated by decoder branches as the input, this module aims to use hand features to enhance the object features to focus more precisely on the interacting region, where $\hat{C}$, $\hat{H}$, and $\hat{W}$ are the channel, height, and width of the feature map.
+Assume we use separate decoder branches for hands and interacting objects, taking the $N^{th}$ layer of hand features $\textbf{D}_{hand}\in\mathbb{R}^{\hat{C}\times\hat{H}\times\hat{W}}$ from the hand decoder branch and the object features $\textbf{D}_{obj}\in\mathbb{R}^{\hat{C}\times\hat{H}\times\hat{W}}$ generated by the object decoder branch as input, this module aims to use hand features to enhance the representation of interacting objects, where $\hat{C}$, $\hat{H}$, and $\hat{W}$ are the channel, height, and width of the feature map.
 First, two stacked convolution operations are used after the LayerNorm layer to generate the query, key, and value.
-The query $Q\in\mathbb{R}^{\hat{C}\times\hat{H}\hat{W}}$ is generated from hand features, and the key $K\in\mathbb{R}^{\hat{C}\times\hat{H}\hat{W}}$ and value $V\in\mathbb{R}^{\hat{C}\times\hat{H}\hat{W}}$ are generated by object features, which can be denoted as follows,
-$Q=f_{conv}(f_{conv}(f_{ln}(\textbf{F}_{hand}))),$ (4)
-$K,V=f_{conv}(f_{conv}(f_{ln}(\textbf{F}_{obj}))).\vspace{-0.3cm}$ (5)
-After that, the query and key are used to perform the hand-guided attention operation, i.e.,
-$\textbf{A}=\delta(Matmul(Norm(Q),Norm(K^{T}))),$ (6)
-$\textbf{F}_{o}=Conv(Matmul(\textbf{A},V)),$ (7)
-where the $\textbf{F}_{o}$ is the output after convolution, the $Matnul$ means the matrix multiplication, and the $\delta(\cdot)$ represents the softmax function.
+The query $Q\in\mathbb{R}^{\hat{C}\times\hat{H}\hat{W}}$ is generated from hand features, and the key $K\in\mathbb{R}^{\hat{C}\times\hat{H}\hat{W}}$ and value $V\in\mathbb{R}^{\hat{C}\times\hat{H}\hat{W}}$ are generated by interacting-object features, which can be denoted as follows:
+$Q=f_{conv}(f_{conv}(f_{ln}(\textbf{D}_{hand}))),$ (7)
+$K,V=f_{conv}(f_{conv}(f_{ln}(\textbf{D}_{obj}))),$ (8)
+where the $f_{conv}(\cdot)$ and $f_{ln}(\cdot)$ represent the convolution and LayerNorm layers. After that, the query, key, and value are used to perform the hand-guided attention operation, i.e.:
+$\textbf{A}=\delta(f_{Matmul}(Norm(Q),Norm(K^{T}))),$ (9)
+$\textbf{F}_{o}=f_{conv}(f_{Matmul}(\textbf{A},V)),$ (10)
+where the $\textbf{F}_{o}$ is the output after convolution, the $f_{Matnul}(x,y)$ means the matrix multiplication between x and y, and the $\delta(\cdot)$ represents the softmax function.
 Then, the input hand features are added to the $\textbf{F}_{o}$ to generate the guided features.
-Finally, the addition between guided features and original object features is performed to enhance the representation of the object features, which can be described as follows,
-$\textbf{F}_{guide}=\textbf{F}_{o}+\textbf{F}_{hand},\quad\hat{\textbf{F}}_{obj}=\textbf{F}_{obj}+\textbf{F}_{guide},$ (8)
-where the $\hat{\textbf{F}}_{obj}$ is the output enhanced object features.
-III-E Object Relation Decoupling Strategy
+Finally, the element-wise addition between guided features and original object features is performed to enhance the representation of the interacting-object features, which can be described as follows:
+$\textbf{F}_{guide}=\textbf{F}_{o}+\textbf{D}_{hand},$ (11)
+${\textbf{D}}_{obj}^{\prime}=\textbf{D}_{obj}+\textbf{F}_{guide},$ (12)
+where the ${\textbf{D}}_{obj}^{\prime}$ is the final enhanced interacting-object features.
+In this way, the relationships between hands and interacting objects are established explicitly by the hand-guided attention mechanism.
+Therefore, hand features can be used as queries to extract more representative and distinguishing object features, improving the interacting-object segmentation performance.
+3.4 Object Relation Decoupling Strategy
 Based on the proposed HOR module, a naive way is to use one decoder branch to predict hands and another decoder branch to predict interacting objects.
-Then, the HOR module can establish the relation between hands and objects, as shown in Sec.
-III-D.
-However, this paper argues that in this way, the three different types of objects, i.e., left-hand objects, right-hand objects, and two-hand objects, are coupled, which will cause confusion and ambiguity when training the network.
-In particular, if the network is trained to predict left-hand objects and right-hand objects based on whether they are in contact with the left and right hands, then two-hand objects that are in contact with both hands simultaneously are likely to be confused with either left-hand objects or right-hand objects.
-Therefore, different from previous work, our method proposes the ORD strategy to decouple the three kinds of objects according to their physical meanings.
-Specifically, we consider two-hand objects to be both left-hand and right-hand objects so that two-hand objects can be represented by the intersection of the other two categories.
-This approach establishes a relationship between two-hand objects and the other two categories that are originally unrelated, enabling the network to concentrate solely on the hand-object interaction without classifying the category.
+Then, the HOR module can establish the relation between hands and interacting objects, as shown in Sec.
+3.3.
+However, this paper argues that in this way, the three different types of interacting objects, i.e., left-hand objects, right-hand objects, and two-hand objects, are coupled.
+To be specific, we observes that the two-hand object interacts with both left and right hands, so it encompasses attributes of both the left-hand and right-hand objects.
+Based on this observation, we contend that two-hand objects can be conceptualized as the intersection of the remaining two categories (left-hand object and right-hand object).
+Therefore, if we remove the two-hand object category during training and represent this category by using the other two categories in inference, the network can be relieved from classifying “whether the object is interacting with both hands at the same time” and focus only on the left- and right-interaction.
+Therefore, different from previous work, our method proposes the ORD strategy to decouple the three kinds of interacting objects according to observed relations.
 In practice, the ORD process can be divided into data preparation, training, and inference.
-First, in the phase of data preparation, we consider the ground truth (GT) of two-hand objects in the dataset as both left-hand objects and right-hand objects, so we generate the new GT of left- and right-hand objects by union operation, as shown in Fig.
-2 (left).
-For example, if the value in the two-hand object GT $\textbf{G}_{to}$ at the position $(i,j)$ is 1, then the value of the left-hand object GT $\textbf{G}_{lo}$ and right-hand object GT $\textbf{G}_{ro}$ at $(i,j)$ will be set to 1.
-The process is shown below,
-$\hat{\textbf{G}}_{lo}=\textbf{G}_{lo}\bigcup\textbf{G}_{to},\quad\hat{\textbf{G}}_{ro}=\textbf{G}_{ro}\bigcup\textbf{G}_{to},$ (9)
-where the $\bigcup$ means the union operation, the $\hat{\textbf{G}}_{lo}$ and $\hat{\textbf{G}}_{ro}$ are the generated GT labels for left-hand objects and right-hand objects.
-Then, in the training process, we only use $\hat{\textbf{G}}_{lo}$ and $\hat{\textbf{G}}_{ro}$ to supervise the two object prediction branches, as shown in Fig. 2, which can be expressed as follows,
-$\hat{\textbf{M}}_{O}^{L}=f_{dec}^{lo}(\textbf{F}_{ehc},\textbf{F},W_{dec}^{lo}),\textbf{F}=\{\textbf{F}_{i}|i=0,1,2\},$ (10)
-$\hat{\textbf{M}}_{O}^{R}=f_{dec}^{ro}(\textbf{F}_{ehc},\textbf{F},W_{dec}^{ro}),\textbf{F}=\{\textbf{F}_{i}|i=0,1,2\},$ (11)
-$\mathscr{L}_{lo}=l_{ce}(\hat{\textbf{M}}_{O}^{L},\hat{\textbf{G}}_{lo}),\quad\mathscr{L}_{ro}=l_{ce}(\hat{\textbf{M}}_{O}^{R},\hat{\textbf{G}}_{ro})),$ (12)
-where the $\hat{\textbf{M}}_{O}^{L}$ and $\hat{\textbf{M}}_{O}^{R}$ are the predicted left-hand objects and right-hand objects.
-The $f_{dec}^{lo}(\cdot)$ and $f_{dec}^{ro}(\cdot)$ represent the decoder branch of left-hand objects and right-hand objects with parameters of $W_{dec}^{lo}$ and $W_{dec}^{ro}$.
-And the $\mathscr{L}_{lo}$ and $\mathscr{L}_{ro}$ denote the loss function of left-hand objects and right-hand objects, which we use the cross entropy $l_{ce}(\cdot)$ in practice.
-In inference, based on the predicted left- and right-hand objects $\hat{\textbf{M}}_{O}^{L}$ and $\hat{\textbf{M}}_{O}^{R}$, the two-hand objects are generated by a simple but effective operation, i.e., intersection, as shown in Fig.
+During data preparation, since two-hand objects encompass attributes of left-hand and right-hand objects, the ground truth (GT) of the two-hand object is treated as the left-hand and right-hand objects to generate new object labels.
+Specifically, as shown in Fig.
+2 (left), the new GT of the left/right-hand object is obtained by the union operation between the original left/right-hand object and two-hand object labels.
+The process is shown below:
+${\textbf{G}}_{lo}^{\prime}=\textbf{G}_{lo}\cup\textbf{G}_{to}=\{(x\in\textbf{G}_{lo})\vee(x\in\textbf{G}_{to})\},$ (13)
+${\textbf{G}}_{ro}^{\prime}=\textbf{G}_{ro}\cup\textbf{G}_{to}=\{(x\in\textbf{G}_{ro})\vee(x\in\textbf{G}_{to})\},$ (14)
+where the $\cup$ means the union operation, the ${\textbf{G}}_{lo}$, ${\textbf{G}}_{ro}$, and ${\textbf{G}}_{to}$ are the original labels for left-hand objects, right-hand objects, and two-hand objects, respectively.
+The ${\textbf{G}}_{lo}^{\prime}$ and ${\textbf{G}}_{ro}^{\prime}$ represent the generated new supervision signals for left-hand objects and right-hand objects.
+Then, during training, we use separate decoder branches to predict the left-hand objects and right-hand objects, as shown in Fig.
+2 (middle).
+The left-hand object decoder output $\textbf{M}_{lo}$ can be obtained by up-sampling the outcome $\textbf{D}_{lo}^{\prime}$ of the HOR module from equation 12, and the right-hand object decoder output $\textbf{M}_{ro}$ can be obtained by up-sampling the outcome $\textbf{D}_{ro}^{\prime}$ of the HOR module from equation 12.
+${\textbf{G}}_{lo}^{\prime}$ and ${\textbf{G}}_{ro}^{\prime}$ are utilized to supervise these two branches, which can be expressed as follows,
+$\mathscr{L}_{lo}=l_{ce}(\textbf{M}_{lo},\textbf{G}_{lo}^{\prime}),$ (15)
+$\mathscr{L}_{ro}=l_{ce}(\textbf{M}_{ro},\textbf{G}_{ro}^{\prime}),$ (16)
+where the $\mathscr{L}_{lo}$ and $\mathscr{L}_{ro}$ denote the loss of left-hand objects and right-hand objects, which we use the cross entropy $l_{ce}(\cdot)$ in practice.
+In inference, after obtaining the decoder output of left- and right-hand objects ${\textbf{M}}_{lo}$ and ${\textbf{M}}_{ro}$, the results of two-hand objects are generated by an ingenious operation, i.e., intersection.
+The final predictions of the left-hand objects and the right-hand objects have to remove the region of the two-hand objects, as shown in Fig.
 2 (right) and follows,
-$\textbf{M}_{O}^{T}=\hat{\textbf{M}}_{O}^{L}\bigcap\hat{\textbf{M}}_{O}^{R},$ (13)
-$\textbf{M}_{O}^{L}=\hat{\textbf{M}}_{O}^{L}-\textbf{M}_{O}^{T},\quad\textbf{M}_{O}^{R}=\hat{\textbf{M}}_{O}^{R}-\textbf{M}_{O}^{T},$
-where the $\bigcap$ means the intersection operation, and the $\textbf{M}_{O}^{L}$ and $\textbf{M}_{O}^{R}$ are the final left- and right-hand object predictions.
-The proposed ORD strategy implicates the physical meaning of the three kinds of objects, especially the meaning of the two-hand objects, i.e., objects that interact with both hands simultaneously.
-Furthermore, in this way, the network can be trained only to predict the objects that interact with the left-/right-hand rather than focusing on classifying the category of the objects, which achieves decoupling the objects and decreasing the confusion and ambiguity of the algorithm.
-III-F Training and inference
-The overall ORMNet is trained by the weighted cross-entropy loss function between predictions and GTs of hands, left-hand objects, right-hand objects, and contact boundaries.
-The overall loss function $\mathscr{L}$ can be computed by,
-$\mathscr{L}=\alpha(l_{ce}(\textbf{M}_{O}^{L},\hat{\textbf{G}}_{lo})+l_{ce}(\textbf{M}_{O}^{R},\hat{\textbf{G}}_{ro}),)+$ (14)
+$\textbf{M}_{O}^{T}={\textbf{M}}_{lo}\cap{\textbf{M}}_{ro}=\{(x\in\textbf{M}_{lo})\wedge(x\in\textbf{M}_{ro})\},$ (17)
+${\textbf{M}}_{O}^{L}={\textbf{M}}_{lo}-\textbf{M}_{O}^{T}=\{(x\in\textbf{M}_{lo})\wedge(x\notin\textbf{M}_{O}^{T})\},$ (18)
+${\textbf{M}}_{O}^{R}={\textbf{M}}_{ro}-\textbf{M}_{O}^{T}=\{(x\in\textbf{M}_{ro})\wedge(x\notin\textbf{M}_{O}^{T})\},$ (19)
+where the $\cap$ means the intersection operation.
+The ${\textbf{M}}_{O}^{L}$, ${\textbf{M}}_{O}^{R}$, and ${\textbf{M}}_{O}^{T}$ are the final left-, right-, and two-hand object predictions, respectively.
+The innovative ORD strategy decouples the two-hand objects during training using pure mathematical intersection and union operations.
+This approach enables the algorithm to prioritize the interaction between hands and objects, thereby diminishing the emphasis on classification and decreasing the confusion and ambiguity of the algorithm.
+3.5 Training and Inference
+The overall ORMNet is trained by the weighted cross-entropy loss function between predictions and GTs of hands, left-hand objects, right-hand objects, and contact boundaries [9].
+The overall loss function $\mathscr{L}$ can be computed by:
+$\mathscr{L}=\alpha(l_{ce}(\textbf{M}_{lo},{\textbf{G}}_{lo}^{\prime})+l_{ce}(\textbf{M}_{ro},{\textbf{G}}_{ro}^{\prime}))+$ (20)
 $\gamma l_{ce}(\textbf{M}_{H},{\textbf{G}}_{h})+\lambda l_{ce}(\textbf{M}_{CB},\textbf{G}_{cb}),$
-where the $l_{ce}(x,y)$ means the cross entropy loss between $x$ and $y$, the $\textbf{M}_{CB}$ and $\textbf{G}_{cb}$ are the prediction and GT of the contact boundary, respectively.
-The GT of the contact boundary is generated by the GT of the hands and interacting objects automatically, which is the overlapping region of the dilated hand and object masks [4].
+where $\alpha,\;\gamma,$ and $\lambda$ denote the loss weight of interacting objects, hands, and contact boundaries [9], respectively.
+$\textbf{M}_{CB}$ presents the output of the contact boundary decoder branch, which is supervised by $\textbf{G}_{cb}$.
+Following previous work [9], the GT of the contact boundary is generated by the overlapping region of the dilated hand and interacting-object masks.
+$\textbf{M}_{H}$ denotes the output of the hand decoder branch, which is supervised by $\textbf{G}_{h}$.
 For inference, as stated in Sec.
-III-E, the network predicts the masks of hands, left-hand objects, and right-hand objects, and then the mask of the two-hand objects can be obtained by the proposed ORD strategy.
-TABLE I: Comparison results on the EgoHOS in-domain test set measured by IoU/Acc and mIoU/mAcc.
+3.4, the decoder branches predict the masks of hands, left-hand objects, and right-hand objects.
+Then, the final predicted masks of left-hand objects, right-hand objects, and two-hand objects can be obtained by the proposed ORD strategy.
+The overall diagram for training and inference is shown in Algorithm 1.
+Algorithm 1 ORMNet
+  ORD strategy - Data Preparation
+        ${\textbf{G}}_{lo}^{\prime}=\textbf{G}_{lo}\bigcup\textbf{G}_{to}$
+        ${\textbf{G}}_{ro}^{\prime}=\textbf{G}_{ro}\bigcup\textbf{G}_{to}$
+  TRAIN
+[…]
+       $\textbf{Prediction: }{\textbf{M}}_{lo},{\textbf{M}}_{ro},\textbf{M}_{H},\textbf{M}_{CB}$
+[…]
+[…]
+       Return: $W$
+  Inference
+       $\textbf{Prediction: }{\textbf{M}}_{lo},{\textbf{M}}_{ro},\textbf{M}_{H},\textbf{M}_{CB}$
+       $\textbf{ORD Strategy - Inference: }\textbf{M}_{O}^{T}={\textbf{M}}_{lo}\bigcap{\textbf{M}}_{ro}$
+                          $\textbf{M}_{O}^{L}={\textbf{M}}_{lo}-\textbf{M}_{O}^{T}$
+                          $\textbf{M}_{O}^{R}={\textbf{M}}_{ro}-\textbf{M}_{O}^{T}$
+       $\textbf{Return: }\textbf{M}_{H},\textbf{M}_{CB},\textbf{M}_{O}^{T},\textbf{M}_{O}^{L},\textbf{M}_{O}^{R}$
+4 Experiments
+4.1 Datasets and Metrics
+To validate the effectiveness of the proposed ORMNet, we compare its performance with different methods on the EgoHOS [9] in-domain test set.
+Additionally, we employed the EgoHOS out-of-domain test set as well as a dataset we developed from the HOI4D [43] dataset, referred to as mini-HOI4D, to assess the generalization capability of our ORMNet.
+EgoHOS.
+The EgoHOS [9] dataset consists of 11,743 egocentric images containing per-pixel segmentation labels of hands and interacting objects gathered from Ego4D [24], EPICKITCHEN [26], THU-READ [44], and their own collected egocentric videos.
+Among them, 8,993 images are used for training, 1,124 images are for validation, 1,126 images are used for in-domain testing, and 500 images are used for out-of-domain testing.
+mini-HOI4D.
+Since the EgoHOS task is an emerging problem, to the best of our knowledge, no suitable datasets other than the EgoHOS dataset can be used to validate the generalization of the model.
+As a result, we selected the relatively similar HOI4D [43] dataset and further processed and aligned its annotations to ensure its applicability for the EgoHOS task requirements.
+Ultimately, we generated a new dataset, named mini-HOI4D, comprising 1,095 egocentric images with corresponding hands and interacting object annotations, which can be used to assess and compare the generalization capabilities of our ORMNet and other approaches.
+This paper uses the most common and widely used metrics in the segmentation task to evaluate the performance, i.e., Intersection-over-Union (IoU) and pixel Accuracy (Acc) for each category.
+Mean IoU (mIoU) and mean accuracy (mAcc) are also employed to indicate the overall performance.
+4.2 Implementation Details
+The experiments were conducted on 4 NVIDIA RTX 6000 Ada GPUs.
+The batch size was set to 12.
+Data preparation included cropping the image to (448, 448) and normalization with the mean of [106.011, 95.400, 87.429], and the standard variation was set to [64.357, 60.889, 61.419].
+Following previous work [9], we also used the Swin Transformer as the backbone.
+The patch size was set to 4, and the window size was set to 12.
+The AdamW optimizer was used in the experiment, the learning rate was set to 1e-5, and the weight decay was set to 0.01.
+The learning rate was set dynamically, i.e., from 0 to 10,000 iterations, the learning rate increased linearly to 1e-4, and from 10,000 iterations to the maximum iteration, the learning rate decreased linearly to 0.
+The maximum number of iterations in the experiment was set to 180,000.
+4.3 Quantitative Results
+4.3.1 In-domain Results
+Table 1: Comparison results on the EgoHOS in-domain test set measured by IoU/Acc and mIoU/mAcc.
+The blue letters represent the best results.
 The ``"`$\prime$"` ` ′ " means the method with data augmentation, the ``"``"`$\diamondsuit$"` ` ♢ " means the method introduced contact boundary, and the ``"``"`$\flat$"` ` ♭ " means the method with data augmentation and contact boundary.
@@
 ORMNet parallel Swin-B 92.34 ||\,|\,|96.64 93.64 ||\,|\,|96.81 60.07 ||\,|\,|71.79 56.69 ||\,|\,|68.71 54.73 ||\,|\,|65.85 71.49 ||\,|\,|79.96
-IV Experiments
-This section first introduces the datasets and metrics we used to validate the EgoHOS performance.
-Second, we compare the proposed ORMNet with state-of-the-art methods in the EgoHOS task.
-Then, we ablate the import design strategy and elements of ORMNet.
-IV-A Datasets and Metrics
-EgoHOS.
-The EgoHOS [4] dataset consists of 11,743 egocentric images containing per-pixel segmentation labels of hands and interacting objects, gathered from Ego4D[55], EPICKITCHEN[56], THU-READ[57], and their own collected egocentric videos.
-Among them, 8,993 images are used for training, 1,124 images are for validation, 1,126 images are used for in-domain testing, and 500 images are used for out-of-domain testing.
-mini-HOI4D.
-HOI4D[58] is a large-scale egocentric image and video dataset with rich annotations for category-level hand-object interaction.
-Frame-wise annotations for action segmentation, motion segmentation, panoptic segmentation, 3D hand pose, and category-level object pose are also provided, together with reconstructed object meshes and scene point cloud.
-In order to test the generalization of the proposed ORMNet, this paper further tests the proposed model with other comparative methods on the HOI4D dataset.
-Since the HOI4D dataset includes the masks of the hand and all objects that have interacted with the hand in a video, we first extract the masks of the objects currently interacting with the hand and delete other objects’ masks.
-Then, we map the label of the interacting objects to the left-hand objects, right-hand objects, and two-hand objects.
-Finally, we sampled a total of 1,095 egocentric images, forming a new dataset mini-HOI4D, to test the generalization of the algorithms on the EgoHOS task.
-Following most related works, this paper uses the most common and widely used metrics in the segmentation task to evaluate the performance, i.e., we use Intersection-over-Union (IoU) and pixel Accuracy (Acc) for each category.
-Mean IoU (mIoU) and mean accuracy (mAcc) are also employed to indicate the overall performance.
-IV-B Implementation Details
-The experiments were conducted on 4 NVIDIA RTX 6000 Ada GPUs.
-The batch size was set to 12.
-Data preparation included cropping the image to (448, 448) and normalization with the mean of [106.011, 95.400, 87.429], and the standard deviation was set to [64.357, 60.889, 61.419].
-The patch size of the Swin transformer encoder was set to 4, and the window size was set to 12.
-The AdamW optimizer was used in the experiment, the learning rate was set to 1e-5, and the weight decay was set to 0.01.
-IV-C Quantitative Results
-1) Comparison results on EgoHOS in-domain test set.
-Table I presents the comparison results of our proposed ORMNet against state-of-the-art methods on the EgoHOS in-domain test set.
-In addition to the parallel (Para.-series) and sequential methods (Seq.-series) proposed in Zhang’s work[4] introduced in detail in Sec.
-II-B, we also select some of the latest transformer-based models with different backbones as parallel methods to compare the effectiveness of the model, e.g., Segformer[13], Segmenter[37], MaskFormer[38], and Mask2former[39], and SCTNet[54].
-All these methods are trained on the EgoHOS train set and tested on the in-domain test set using the same parameter setting as the proposed ORMNet.
+Comparison results on EgoHOS in-domain test set.
+To validate the effectiveness of the proposed ORMNet, we conducted experiments by training on the EgoHOS train set and testing on the EgoHOS in-domain test set.
+Due to the novelty of the EgoHOS task, we selected a set of novel state-of-the-art general segmentation algorithms [16, 46, 48, 49, 45] as compared parallel methods.
+Furthermore, we ensured a comprehensive comparison by incorporating all existing methods [9] tailored for the EgoHOS task in this evaluation.
+The results are presented in Table 1.
 In general, the proposed ORMNet far outperforms all previous methods in most categories on both metrics.
-Specifically, compared with the previous best parallel method in the $8^{th}$ row, Mask2Former, our method has demonstrated pronounced advancements across various categories and exhibited particularly notable enhancements in object segmentation.
+Specifically, compared with the previous best parallel method - Mask2Former [49], our method has demonstrated pronounced advancements across various categories and exhibited particularly notable enhancements in interacting-object segmentation.
 For example, our method showcases substantial enhancements of 15.85%, 10.64%, and 3.6% on left-, right-, and two-hand objects, respectively.
-Furthermore, compared with the results of the previous best sequential methods $\text{Seq}^{\flat}$, our ORMNet improves IoU by more than 4% on both the left and right hands.
+Moreover, compared with the results of the previous best sequential methods $\text{Seq}^{\flat}$, our ORMNet improves IoU by more than 4% on both the left and right hands.
 The most substantial enhancement is observed in the right-hand objects, surpassing 12%.
 Finally, compared with the previous best method, our proposed method attains a commendable advancement of 4.33% and 4.55% in mIoU and mAcc, respectively.
-This is because we propose an ORD strategy to decouple the relationship between object categories, which reduces the burden of network learning.
-Furthermore, we propose the HOR module to strengthen the correlation between hand features and object features, significantly improving the segmentation accuracy of hands and objects.
-TABLE II: Comparison results on the egohos out-of-domain test set.
+These results profoundly demonstrate the state-of-the-art segmentation performance achieved by the proposed ORMNet on the basis of end-to-end optimization, leveraging the HOR module and the ORD strategy to model the correlations between hands and objects as well as objects and objects.
+Table 2: Comparison results on the EgoHOS out-of-domain test set. The blue letters represent the best results. The ``"``"`$\flat$"` ` ♭ " means the method with data augmentation and contact boundary.
 Method Type Backbone Left hand Right hand Left-hand objects Right-hand objects Two-hand objects Overall
@@
 ORMNet parallel Swin-B 94.47 ||\,|\,|97.09 94.41 ||\,|\,|96.69 51.56 ||\,|\,|72.30 36.80 ||\,|\,|60.90 41.84 ||\,|\,|46.28 63.82 ||\,|\,|74.65
-TABLE III: Comparison results on the mini-HOI4D test set. Note: The left-hand objects are not considered to compare due to few samples.
+4.3.2 Out-of-domain Results
+1) Comparison results on EgoHOS out-of-domain test set.
+In order to test the generalization performance of different models, the EgoHOS dataset provides 500 out-of-domain images, which are captured from YouTube and annotated manually.
+Thus, We compare ORMNet with different parallel and sequential models by directly testing saved best models on the EgoHOS out-of-domain test set, results are shown in Table 2.
+It can be observed that the results of our proposed ORMNet on the two metrics of most categories have obtained significant generalization performance.
+Among them, the IoU of the left and right hand exceed 94%, and the Acc.
+of these two categories exceed 96%.
+The segmentation results on various interacting-object categories also achieve balanced performance.
+Finally, our ORMNet achieves state-of-the-art generalization performance in overall mIoU and mAcc metrics on the EgoHOS out-of-domain test set.
+Table 3: Comparison results on the mini-HOI4D test set. The blue letters represent the best results. The ``"``"`$\flat$"` ` ♭ " means the method with data augmentation and contact boundary.
 Method Type Backbone Left hand Right hand Right-hand objects Two-hand objects Overall
@@
 ORMNet parallel Swin-B 70.39 ||\,|\,|97.79 89.76 ||\,|\,|93.09 27.56 ||\,|\,|49.35 60.08 ||\,|\,|68.01 61.95 ||\,|\,|77.06
-2) Comparison results on EgoHOS out-of-domain test set.
-In order to test the generalization performance of different models, the EgoHOS dataset provides 500 out-of-domain images, which are captured from YouTube and annotated manually.
-Table II illustrates the comparison results of our method with the previous methods on the EgoHOS out-of-domain test set measured by IoU and Acc.
-Please note that some of the results in the table are from direct inference using the officially released best model (if available), and other results are from direct inference on the best model trained on our own.
-It can be observed that the results of our proposed ORMNet on the two metrics of most categories have obtained competitive generalization performance.
-Among them, the IoU of the left and right hand exceed 94%, and the Acc.
-of these two categories exceed 96%.
-The segmentation results on various object categories also achieve balanced performance.
-Finally, our ORMNet achieves state-of-the-art performance in overall mIoU and mAcc metrics and demonstrates excellent generalization performance on the out-of-domain test set.
-3) Comparison results on mini-HOI4D dataset.
-In addition to the evaluation on the EgoHOS out-of-domain test set, we also assess the model’s generalization capabilities using a completely different mini-HOI4D dataset.
-It should be noted that due to the limited samples of left-hand objects in the original HOI4D[58] dataset, the generated mini-HOI4D dataset does not include samples of left-hand objects.
-Consequently, no testing is performed on left-hand objects.
-The comparison results on the mini-HOI4D data set are shown in Table III.
-Similarly to the EgoHOS out-of-domain test set, the results on the mini-HOI4D dataset are directly inferred by utilizing trained and saved checkpoints without further training procedures.
+2) Comparison results on mini-HOI4D dataset.
+In order to comprehensively validate the generalizability of the proposed method, we also assess the generalization capabilities of algorithms using a completely different mini-HOI4D dataset.
+It should be noted that since the original HOI4D [43] dataset contains very few samples of left-hand objects, we do not perform testing on left-hand object category.
+We also utilized the saved best models to test on the mini-HOI4D dataset without further training procedures, the comparison results are shown in Table 3.
 Overall, our proposed ORMNet outperforms all the previous parallel and sequential methods concerning mIoU and mAcc.
-Our method excels notably in hand segmentation, i.e., 1.57% and 3.71% improvements on IoU and Acc for the right hand compared with the $8^{th}$ row.
-Our method also competes well in interacting object segmentation, yielding commendable outcomes.
+Our method excels notably in hand segmentation, i.e., 1.57% and 3.71% improvements on IoU and Acc for the right hand compared with the previous best Mask2Former model.
+Our method also competes well in interacting-object segmentation, yielding commendable outcomes.
 Therefore, it can be deduced from the results of the EgoHOS out-of-domain and mini-HOI4D test sets that our proposed ORMNet exhibits superior generalization capabilities in contrast to the comparative methods.
-[…]
-TABLE IV: Ablation results on the EgoHos in-domain test set under the same setting measured by IoU and mIoU.
+Table 4: Ablation results on the EgoHOS in-domain test set under the same setting measured by IoU and mIoU.
 Lh: left hand, rh: right hand, lo: left-hand object, ro: right-hand object, to: two-hand object.
@@
 Basic+ORD+HOR 92.34 93.64 60.07 56.69 54.73 71.49
-TABLE V: The experimental results using different hyper-parameters.
-$\alpha$, $\gamma$, and $\lambda$ are the weights of objects, hands, and CB losses.
-Lh: left hand, rh: right hand, lo: left-hand object, ro: right-hand object, to: two-hand object.
+4.3.3 Ablation Study
+In order to verify the effectiveness of the proposed method alongside its corresponding sub-module and strategy, we conduct the ablation study in this section.
+The experimental results are shown in Table 4.
+The basic model uses Swin-B as the encoder and uses three decoder branches to predict hands, contact boundaries, and interacting objects.
+This basic model is a naive solution for the EgoHOS task and does not model any relationships between hands and interacting objects, so its results are relatively poor.
+The $2^{nd}$ and $3^{rd}$ rows of the table show the results after adding the ORD strategy and HOR module independently, which proves that adding them to the basic model is beneficial.
+The last row is the result after integrating the ORD strategy and the HOR module.
+By explicitly modeling the relations between hands and objects as well as objects and objects, the segmentation accuracy achieves its peak performance.
+Generally speaking, although the effect improvement brought by the ORD strategy is higher than that of the HOR module, using these two modules and strategies simultaneously can further improve the segmentation performance, which verifies the effectiveness of our proposed method and sub-methods.
+4.3.4 Hyper-parameter Study
+The method presented in this paper requires setting multiple hyper-parameters for total training loss, i.e., $\alpha,\gamma$, and $\lambda$.
+This section conducts experiments on the configurations of different hyper-parameters, results are shown in Table 5.
+We assign the values of $\alpha$, $\gamma$, and $\lambda$ based on the relative significance of the interacting objects, hands, and contact boundaries.
+In the $1^{st}$ row, all categories are considered equal importance, while in subsequent experiments (i.e., $2^{nd},3^{rd}$ and $4^{th}$), we regard the interacting objects, CBs, and hands as the most crucial, respectively.
+The results lead to the inference that amplifying the importance of hands can yield enhancements in the method’s performance, which can be explained intuitively because the proposed HOR module leverages hands as prior knowledge for interacting-object feature enhancement.
+Table 5: The experimental results using different hyper-parameters. $\alpha$, $\gamma$, and $\lambda$ are the weights of interacting objects, hands, and CB losses.
 $\alpha$ $\gamma$ $\lambda$     Lh (%) Rh (%) Lo (%) Ro (%) To (%)     mIoU (%)
@@
 0.5 1.0 0.5     92.34 93.64 60.07 56.69 54.73     71.49
-4) Ablation study.
-In order to verify the effectiveness of the proposed method alongside its corresponding sub-module and strategy, we conduct the ablation study in this section.
-The experimental results are shown in Table IV.
-The basic model uses Swin-B as the encoder and uses three decoder branches to predict hands, contact boundaries, and objects.
-This basic model is a naive solution for the EgoHOS task and does not model any relationships between hands and interacting objects, so its results are relatively poor.
-The $2^{nd}$ and $3^{rd}$ rows of the table show the results after adding the ORD strategy and HOR module independently, which proves that adding them to the basic model is beneficial.
-The last row is the result after integrating the ORD strategy and the HOR module.
-By explicitly modeling both the hand-object relationship and the object-object relationship, the segmentation accuracy achieves its peak performance.
-Generally speaking, although the effect improvement brought by the ORD strategy is higher than that of the HOR module, using these two modules and strategies simultaneously can further improve the segmentation performance, which verifies the effectiveness of our proposed method and sub-methods.
-Figure 5: Visualization results of the ORMNet compared with the sequential method on the EgoHOS out-of-domain test set (left) and mini-HOI4D dataset (right).
-Red: left hand, dark blue: right hand, pink: left-hand objects, light blue: right-hand objects, and green: two-hand objects.
-5) Hyper-parameter experiments.
-While identifying the optimal hyper-parameters is not the primary focus of this paper, this section still conducts experiments on the configurations of different hyper-parameters.
-The experimental results are shown in Table V.
-We assign the value of $\alpha$, $\gamma$, and $\lambda$ based on the relative significance of the interacting objects, hands, and contact boundaries.
-In the $1^{st}$ row, all categories are considered equal importance, while in subsequent experiments (i.e., $2^{nd},3^{rd}$ and $4^{th}$), we regard the objects, CBs, and hands as the most crucial, respectively.
-The results lead to the inference that amplifying the importance of hands can yield enhancements in the method’s performance, which can be explained intuitively because the proposed HOR module leverages hands as prior knowledge for feature enhancement.
-IV-D Qualitative Results
+4.4 Qualitative Results
+4.4.1 In-domain Results
 […]
-2) Results on EgoHOS out-of-domain test set.
-We directly used the saved checkpoint to conduct experiments on the EgoHOS out-of-domain test set and compared it with the $\text{Seq}^{\flat}$ method.
-The visualized experimental results are shown in Fig.
+Figure 5: Visualization results of two-hand objects on the EgoHOS in-domain test set compared with sequential method $\text{Seq}^{\flat}$ [9].
+[…]
+Figure 6: Visualization results of the ORMNet compared with the sequential method [9] on the EgoHOS out-of-domain test set (left) and mini-HOI4D dataset (right).
+The main improvements are highlighted in the dashed yellow box.
+4.4.2 Out-of-domain Results
+1) Results on EgoHOS out-of-domain test set.
+We visualize the comparative results on the EgoHOS out-of-domain test set with the $\text{Seq}^{\flat}$ method in Fig.
 6 (left).
-According to the visualization results, our ORMNet has stronger generalization than the compared sequential model.
+The figure shows that our ORMNet has stronger generalization than the compared sequential model.
 Specifically, our method can segment and identify left and right hands accurately, as shown in the $3^{rd}$ to the last rows of results.
-Furthermore, our method can precisely segment and recognize different kinds of objects, ensuring consistent segmentation of interactive objects.
-3) Results on mini-HOI4D dataset.
-The experiments on the mini-HOI4D dataset are also conducted directly using the saved checkpoints.
-The experimental results are shown in Fig.
+Furthermore, our method can precisely segment and recognize different kinds of interacting objects, as shown in the first and second rows in the figure.
+2) Results on mini-HOI4D dataset.
+We visualize the comparative results on the mini-HOI4D test set with the $\text{Seq}^{\flat}$ method in Fig.
 6 (right).
-It can be observed from the experimental results that the contrasting sequential method easily confuses the left and right hands, such as $1^{st}$, $2^{nd}$, $4^{th}$, and $5^{th}$ rows of the results.
+It can be observed that the contrasting sequential method easily confuses the left and right hands, such as $1^{st}$, $2^{nd}$, $4^{th}$, and $5^{th}$ rows of the results.
 In contrast, our method exhibits robust generalization capabilities coupled with outstanding segmentation performance.
-In addition, although some tiny parts are not considered, such as the handle of the bucket in the third row, our method also significantly excels in segmenting two-hand objects.
-V Conclusion
-This paper introduces an end-to-end high-performance Object-centric Relationship Modeling Network to effectively handle the egocentric hand-object segmentation task.
-First, the HOR module is proposed to facilitate the modeling of hand-object interactions through hand-guided cross-attention, enhancing the network’s adaptability to learn precise representations.
-Second, this paper designs the ORD strategy to decouple the two-hand objects, thereby reducing the network’s confusion about object segmentation.
-Experimental results demonstrate that the proposed ORMNet exhibits SOTA segmentation efficacy and robust generalization capabilities across varied datasets.
-VI Acknowledgements
+5 Conclusion
+This paper introduces the ORMNet to effectively handle the EgoHOS task while achieving end-to-end optimization by incorporating two parts.
+First, the HOR module is proposed to facilitate the modeling of hand-object interactions through hand-guided cross-attention, enhancing the network’s adaptability to learn comprehensive interacting-object representations.
+Second, this paper designs an ingenious ORD strategy to decouple the two-hand objects, thereby reducing the confusion of the network about interacting-object classification and emphasizing the left/right interactions.
+However, there remains a limitation that needs to be addressed.
+Specifically, the EgoHOS task focuses solely on predicting hands (left- and right-hand) and interacting objects (left-hand object, right-hand object, and two-hand object) without inferring the object’s name or semantic information.
+Consequently, this limits the comprehensive description of people’s actions.
+To address this issue, we plan to enhance and refine the EgoHOS task in future work.
+Our goal is to incorporate the prediction of interacting-object semantic information alongside hand-object segmentation, thereby expanding the practical applicability of the algorithm in real-world scenarios.
+6 Acknowledgements
 The research work was conducted in the JC STEM Lab of Machine Learning and Computer Vision funded by The Hong Kong Jockey Club Charities Trust.